### v1

# Emulation of Complex Matrix Multiplication based on the Chinese Remainder Theorem

Abstract
[…]
I Introduction
I-A Background
General-purpose computing architectures have increasingly adopted low-precision matrix multiplication units in Graphics Processing Units (GPUs), Tensor Processing Units (TPUs), and Neural Processing Units (NPUs). Representative examples include Tensor Cores in NVIDIA GPUs [tensorcore], Matrix Cores in AMD GPUs [matrixcore], the MXU in Google TPUs [tpu], and the MAC array in Intel NPUs [intelXecore]. These units offer substantially higher arithmetic throughput than high-precision units and have become central components of modern accelerator design. Table I shows the theoretical throughput performance in TFLOPS or TOPS of AMD [MI300A, MI300X, MI325X, MI350X, MI355X] and NVIDIA GPUs [A100, H100, Blackwell, RTX5080]. In recent years, such units have also begun to appear as standard features even in laptop-class processors [copilot+pc, SME], further broadening their availability. While these units were originally motivated by machine-learning workloads, their architectural characteristics also present an opportunity to accelerate high-precision computations by emulating them on low-precision hardware. Consequently, emulation-based high-precision matrix multiplication has emerged as an active research topic in the high-performance computing community [brower2025mixedprecisionabinitiotensor, abdelfattah2025analysisfloatingpointmatrixmultiplication, schwarz2025guaranteed, mukunoki2025sparseiterativesolversusing, dongarra2024hardwaretrendsimpactingfloatingpoint, 11196192, 11196413].
TABLE I: TFLOPS and TOPS of AMD and NVIDIA GPUs [MI300A, MI300X, MI325X, MI350X, MI355X, A100, H100, Blackwell, RTX5080]
GPU FP64 FP32 FP16 FP8 INT8
Instinct MI300A CDNA3 122.6 122.6 980.6 1961.2 1961.2
Instinct MI300X CDNA3 163.4 163.4 1307.4 2614.9 2614.9
Instinct MI325X CDNA3 163.4 163.4 1307.4 2614.9 2614.9
Instinct MI350X CDNA4 72.1 144.2 2306.9 4613.7 4614
Instinct MI355X CDNA4 78.6 157.3 2516.6 5033.2 5033.2
Ampere A100 SXM 19.5 19.5 312 624 624
Hopper H100 SXM 66.9 66.9 989.4 1978.9 1978.9
Blackwell B200 SXM 37 75 2250 4500 4500
Blackwell Ultra B300 SXM 1.2 75 2250 4500 150
GeForce RTX 5080 Blackwell 0.88 56.3 112.6 450.2 450.2
I-B Contributions
A prior line of work introduced the Ozaki-I [ozaki2012error, ozaki2013generalization, mukunoki2020, mukunoki2025dgemmfp64arithmetic, ootomo2024dgemm, uchino2025Performance] and Ozaki-II [ozaki-scheme2, uchino_ozaki2] schemes as frameworks for emulating high-precision matrix multiplication using low-precision matrix engines, demonstrating high performance and power efficiency on contemporary GPUs. In parallel, the Ozaki-I scheme has recently been incorporated into cuBLAS (from version 13.0 Update 2), and this implementation supports complex matrix multiplication. However, for real-valued matrix multiplication, previous studies have reported that the Ozaki-II scheme achieves higher performance than Ozaki-I. This suggests that a complex extension of the Ozaki-II scheme would likewise outperform the complex Ozaki-I approach. Despite this expectation, no complex-valued formulation of Ozaki-II has been developed to date. This absence is notable because complex matrix multiplication is a key computational kernel in many HPC applications, and its performance can dominate the overall runtime of large-scale scientific simulations.
To address this gap, the present work extends the Ozaki-II scheme to single- and double-precision complex matrix multiplication (CGEMM and ZGEMM, respectively). We develop emulation algorithms for CGEMM and ZGEMM by combining the Ozaki-II scheme with a Karatsuba-based formulation of complex arithmetic. The resulting methods maintain numerical accuracy comparable to standard CGEMM and ZGEMM while effectively exploiting the high throughput of low-precision matrix engines. In addition, several improvements applicable to existing SGEMM and DGEMM emulation techniques are introduced, further enhancing the efficiency of the Ozaki schemes in the real domain. The proposed algorithms are released as a portable library supporting both CUDA and HIP backends, enabling execution on NVIDIA GPUs and AMD GPUs. Performance is evaluated on an NVIDIA GH200 Grace Hopper Superchip, an NVIDIA B200 Blackwell GPU, an NVIDIA GeForce RTX 5080 Blackwell GPU, and an AMD Instinct MI300X GPU, and compared against cuBLAS and hipBLAS implementations. For sufficiently large matrix sizes, the proposed methods surpass both vendor-optimized libraries and, in particular, achieve higher throughput than the cuBLAS Ozaki-I–based ZGEMM emulation, demonstrating the performance potential of Ozaki-II-based emulation for CGEMM and ZGEMM. We also provide performance models for the proposed methods, and the predicted throughput closely aligns with the measured results across all tested configurations.
The remainder of this paper is organized as follows. Section II introduces the Ozaki-II scheme for real-valued matrix multiplication. Section III presents the proposed efficient methods for emulating CGEMM and ZGEMM. Section IV shows the numerical results comparing the proposed emulation with the vender-optimized GEMM routines. Section V presents the concluding remarks. Note that in this paper, we do not consider matrix shapes for which performance becomes memory-bound, such as small matrices or highly skinny matrices, because low-precision matrix engines do not achieve high throughput on such workloads.
II Ozaki-II Scheme for Real Matrix Multiplication
The Ozaki-II scheme is an emulation method of matrix multiplication based on a large-integer multiplication technique using the Chinese Remainder Theorem (CRT). The scheme for matrix multiplication of $A\in\mathbb{R}^{m\times k}$ and $B\in\mathbb{R}^{k\times n}$ consists of the following three steps:
  1. 1.
    Convert $A$ and $B$ to $A^{\prime}\in\mathbb{Z}^{m\times k}$ and $B^{\prime}\in\mathbb{Z}^{k\times n}$, respectively, by applying diagonal scaling and truncation.
  2. 2.
    Compute the integer matrix product $C^{\prime}\approx A^{\prime}B^{\prime}$ based on the CRT.
  3. 3.
    Convert $C^{\prime}$ back to floating-point format $C\in\mathbb{R}^{m\times n}$ by inversely scaling.
The integer matrix multiplication in step 2 is performed by decomposing the elements of $A^{\prime}$ and $B^{\prime}$ into residues with respect to a predefined set of pairwise coprime moduli. Let $p_{1},p_{2},\dots,p_{N}\in\mathbb{N}$ denote these moduli and define $P:=\prod_{\ell=1}^{N}p_{\ell}$. Let $\mathrm{mod}(x,p_{\ell})$ for $x\in\mathbb{Z}$ denote the symmetric remainder of $x/p_{\ell}$, i.e., $\mathrm{mod}(x,p_{\ell})=x-p_{\ell}\cdot\mathrm{round}(x/p_{\ell})$, where $\mathrm{round}(\cdot)$ rounds its input to the nearest integer. Each element of $A^{\prime}$ and $B^{\prime}$ is mapped to its residue representation modulo each $p_{\ell}$ as
$(A^{\prime}_{\ell})_{ij}:=\bmod(a^{\prime}_{ij},p_{\ell}),\quad(B^{\prime}_{\ell})_{ij}:=\bmod(b^{\prime}_{ij},p_{\ell}).$ (1)
Then, the corresponding matrix products are computed independently for each $p_{\ell}$ as
$C^{\prime}_{\ell}:=A^{\prime}_{\ell}B^{\prime}_{\ell}.$ (2)
This procedure yields a set of partial results
$C^{\prime}_{\ell}\equiv A^{\prime}B^{\prime}\pmod{p_{\ell}},\quad\ell=1,\dots,N.$ (3)
The final integer matrix $C^{\prime}$ is then reconstructed from $C^{\prime}_{\ell}$ by applying the CRT as
$C^{\prime}:=\sum_{\ell=1}^{N}\frac{P}{p_{\ell}}q_{\ell}\cdot C^{\prime}_{\ell}.$
Then, we obtain
$C^{\prime}\equiv A^{\prime}B^{\prime}\pmod{P}$,
namely, $c^{\prime}_{ij}=\mathrm{mod}\left((A^{\prime}B^{\prime})_{ij},P\right)+zP$ for $z\in\mathbb{Z}$. If the condition
$2\sum_{h=1}^{k}|a^{\prime}_{ih}||b^{\prime}_{hj}|<P$ (4)
is satisfied for all $(i,j)$ pairs, the possible value of $c^{\prime}_{ij}$ is only $\mathrm{mod}\left((A^{\prime}B^{\prime})_{ij},P\right)$ and the CRT guarantees that the reconstructed $C^{\prime}$ is unique. Therefore, the scaling vector used in step 1 must be chosen such that the bound (4) is ensured for $A^{\prime}$ and $B^{\prime}$. Increasing the number of moduli $p_{\ell}$ enlarges $P$, thereby relaxing the uniqueness condition. However, this also increases the number of matrix multiplications required, leading to higher computational cost.
In previous work [uchino_ozaki2], an algorithm was proposed in which each modular matrix multiplication (3) is executed using INT8 matrix engines. Algorithm 1 shows the outline. Hereafter, we assume that $k\leq 2^{17}$ to perform INT8 matrix multiplication without error. In [uchino_ozaki2], step iv at the line 7 of Algorithm 1 uses the standard (non-symmetric) modulo operation that returns values in $0,\,p_{\ell})$ and $E^{\prime}_{\ell}$ are UINT8 matrices. In DGEMM emulation, step v at the line 7 evaluates $S=\sum_{\ell=1}^{N}P/p_{\ell}\cdot q_{\ell}\cdot E^{\prime}_{\ell}$ by the unevaluated sum $S_{1}+S_{2}$ as
$S\approx S_{1}+S_{2},\quad S_{1}:=\sum_{\ell=1}^{N}s_{\ell 1}E^{\prime}_{\ell},\quad S_{2}:=\sum_{\ell=1}^{N}s_{\ell 2}E^{\prime}_{\ell},$ (5)
where $s_{\ell 1}$ contains the upper $(53-8-\lceil\log_{2}N\rceil)$ bits of $P/p_{\ell}\cdot q_{\ell}$, $s_{\ell 2}$ stores the remaining lower bits rounded to double precision, and $s_{\ell 1}+s_{\ell 2}\approx P/p_{\ell}\cdot q_{\ell}$. In (5), the value of $s_{\ell 1}$ is limited by the precision of the double-precision significand. Since each element of $E^{\prime}_{\ell}$ is an 8-bit unsigned integer, the product $s_{\ell 1}E^{\prime}_{\ell}$ requires 8 bits of significand to represent the contribution of $E^{\prime}_{\ell}$. Moreover, accumulating $N$ such terms requires an additional $\lceil\log_{2}N\rceil$ bits to avoid overlap between partial sums. Thus, the high-order part $s_{\ell 1}$ may occupy at most $(53-8-\lceil\log_{2}N\rceil)$ bits of precision. This bit allocation ensures that all products $s_{\ell 1}E^{\prime}_{\ell}$ can be accumulated in $S_{1}$ without rounding error. In contrast, the present work adopts a symmetric modulo operation in step iv at the line 7, ensuring that each $E^{\prime}_{\ell}$ becomes INT8 from UINT8. Consequently, the upper part $s_{\ell 1}$ of $P/p_{\ell}\cdot q_{\ell}$ can use $(53-7-\lceil\log_{2}N\rceil)$ bits instead of $(53-8-\lceil\log_{2}N\rceil)$ bits, gaining one additional effective bit compared with the previous method. This slight but meaningful increase in representable precision improves the numerical robustness of the reconstruction step without changing the overall algorithmic structure.
Algorithm 1 Outline of Ozaki-II scheme using INT8 matrix engines for real matrix multiplication [uchino_ozaki2]. The function $\mathrm{trunc}(\cdot)$ returns the integer part of its input.
[…]
3:Determine the moduli $p_{\ell}\leq 256$ for $\ell=1,\dots,N$.
4:Determine modular multiplicative inverses $q_{\ell}$ of $P/p_{\ell}$ modulo $p_{\ell}$ for $\ell=1,\dots,N$.
5:Determine the scaling vectors $\mu\in\mathbb{R}^{m}$ and $\nu\in\mathbb{R}^{n}$.
6:Convert AA and $B$ to $A^{\prime}\in\mathbb{Z}^{m\times k}$ and $B^{\prime}\in\mathbb{Z}^{k\times n}$:
  1. i)
    $a^{\prime}_{ij}:=\mathrm{trunc}(a_{ij}\cdot\mu_{i})$
  2. ii)
    $b^{\prime}_{ij}:=\mathrm{trunc}(b_{ij}\cdot\nu_{j})$
7: Compute $C^{\prime}\approx A^{\prime}B^{\prime}$ based on the CRT:
  1. i)
    $(A^{\prime}_{\ell})_{ij}:=\bmod(a^{\prime}_{ij},p_{\ell})$ for $\ell=1,\dots,N$
  2. ii)
    $(B^{\prime}_{\ell})_{ij}:=\bmod(b^{\prime}_{ij},p_{\ell})$ for $\ell=1,\dots,N$
  3. iii)
    $D^{\prime}_{\ell}:=A^{\prime}_{\ell}B^{\prime}_{\ell}$ using INT8-GEMM for $\ell=1,\dots,N$
  4. iv)
    $(E^{\prime}_{\ell})_{ij}:=\bmod(d^{\prime}_{ij},p_{\ell})$ for $\ell=1,\dots,N$
  5. v)
    $S:=\sum_{\ell=1}^{N}P/p_{\ell}\cdot q_{\ell}\cdot E^{\prime}_{\ell}$
  6. vi)
    $c^{\prime}_{ij}:=\bmod(s_{ij},P)$
8:Convert $C^{\prime}$ back to floating-point format $C\in\mathbb{R}^{m\times n}$:
  1. i)
    $C:=\mathrm{diag}(\mu)^{-1}C^{\prime}\mathrm{diag}(\nu)^{-1}$
To satisfy the condition (4) efficiently, the scaling vectors used in step 1 must be determined appropriately. Previous work introduced two strategies for computing this scaling: a fast mode and an accurate mode. The fast mode estimates an upper bound of $\sum_{h=1}^{k}|a^{\prime}_{ih}||b^{\prime}_{hj}|$ via the Cauchy–Schwarz inequality and determines the scaling vector analytically based on this estimate. This approach incurs minimal preprocessing cost, but the bound obtained from the Cauchy–Schwarz inequality tends to be overly conservative. As a consequence, a larger $P$ (i.e., more moduli) is often required to satisfy the uniqueness condition, increasing the computational workload in step 2. In contrast, the accurate mode estimates the upper bound by performing an auxiliary matrix multiplication using INT8 matrix engines. Although this mode requires additional computation during step 1, the direct estimation yields a much tighter bound than that obtained via the Cauchy–Schwarz inequality. This typically enables the use of fewer moduli, reducing the number of modular matrix multiplications and potentially improving overall performance.
III Proposed Methods
This section presents the extension of the Ozaki-II scheme to complex matrix multiplication, targeting high-performance execution on INT8 matrix engines. Let $A=A_{R}+iA_{I}\in\mathbb{C}^{m\times k}$ for $A_{R},A_{I}\in\mathbb{R}^{m\times k}$ and $B=B_{R}+iB_{I}\in\mathbb{C}^{k\times n}$ for $B_{R},B_{I}\in\mathbb{R}^{k\times n}$, and we compute $C=C_{R}+iC_{I}:=AB$, where $C_{R},C_{I}\in\mathbb{R}^{m\times n}$. A standard approach to handle complex matrix multiplication is to employ the following representation:
$\hat{M}:=\begin{pmatrix}M_{R}&-M_{I}\\ M_{I}&M_{R}\end{pmatrix}$ (6)
for $M\in\{A,B,C\}$. This representation forms the basis for both steps 1 and 2 in the complex Ozaki-II scheme.
III-A INT8 Complex Matrix Multiplication
Using (6), $C_{R}$ and $C_{I}$ are obtained by
$\begin{pmatrix}C_{R}\\ C_{I}\end{pmatrix}=\begin{pmatrix}A_{R}&-A_{I}\\ A_{I}&A_{R}\end{pmatrix}\begin{pmatrix}B_{R}\\ B_{I}\end{pmatrix}$ (7)
or
$\begin{pmatrix}C_{I}&C_{R}\end{pmatrix}=\begin{pmatrix}A_{I}&A_{R}\end{pmatrix}\begin{pmatrix}B_{R}&-B_{I}\\ B_{I}&B_{R}\end{pmatrix}$. (8)
The formulations in (7) and (8) allow complex matrix multiplication to be computed using a single real GEMM applied to an expanded matrix. For the formulations, the applicability condition on the inner dimension changes from $k\leq 2^{17}$ to $k\leq 2^{16}$. Another standard approach to handle complex matrix multiplication is the Karatsuba algorithm. The Karatsuba algorithm expresses the complex product using three real matrix multiplications as
$C_{R}:=D-E,\quad C_{I}:=F-D-E$ (9)
for
$\begin{cases}D:=A_{R}B_{R},\\ E:=A_{I}B_{I},\\ F:=(A_{R}+A_{I})(B_{R}+B_{I}),\end{cases}$ (10)
[…]
Figure 1: Performance comparison of four INT8-based GEMM strategies on NVIDIA H100 NVL with CUDA Toolkit 12.8.61. The first two methods use a single INT8 matrix multiplication with dimensions $(2h,h,2h)$ and $(h,2h,2h)$, respectively. The third method performs three INT8 multiplications of size $(h,h,h)$, while the fourth applies the same three-multiplication scheme but introduces blocking along the $n$ dimension.
III-B Scaling Vector Determination
First, we convert $\hat{A}$ and $\hat{B}$ to integer matrices $\hat{A}^{\prime}$ and $\hat{B}^{\prime}$ as
$\hat{a}^{\prime}_{ij}:=\mathrm{trunc}(\hat{a}_{ij}\cdot\mu_{i}),\quad\hat{b}^{\prime}_{ij}:=\mathrm{trunc}(\hat{b}_{ij}\cdot\nu_{j}).$
We also consider two strategies: a fast mode and an accurate mode. From $|\hat{a}^{\prime}_{ij}|\leq|\hat{a}_{ij}|\cdot\mu_{i}$ and $|\hat{b}^{\prime}_{ij}|\leq|\hat{b}_{ij}|\cdot\nu_{j}$,
$\sum_{h=1}^{k}|\hat{a}^{\prime}_{ih}||\hat{b}^{\prime}_{hj}|\leq\mu_{i}\nu_{j}\cdot\sum_{h=1}^{k}|\hat{a}_{ih}||\hat{b}_{hj}|.$
In the fast mode, using the Cauchy–Schwarz inequality yields
$\sum_{h=1}^{k}|\hat{a}_{ih}||\hat{b}_{hj}|\leq\|\hat{a}_{i,:}\|_{2}\|\hat{b}_{:,j}\|_{2}.$
From (6), the row norms of $\hat{A}$ satisfy $\|\hat{a}_{i,:}\|_{2}=\|\hat{a}_{i+m,:}\|_{2}$ for $1\leq i\leq m$, and the column norms of $\hat{B}$ satisfy $\|\hat{b}_{:,j}\|_{2}=\|\hat{b}_{:,j+n}\|_{2}$ for $1\leq j\leq n$. Thus, the scaling values of rows $i$ and $i+m$ of $\hat{A}$ are identical, and the same holds for columns $j$ and $j+n$ of $\hat{B}$. Consequently, the scaling vectors reduce to $\mu\in\mathbb{R}^{m}$ and $\nu\in\mathbb{R}^{n}$. The elements of $\mu$ and $\nu$ can then be determined in the same manner as in the real-valued Ozaki-II scheme [uchino_ozaki2], i.e.,
$\mu_{i}$ $\leq 2^{\lfloor\frac{\log_{2}(P-1)-1}{2}-1-\max\left(1,\ 0.5\cdot\log_{2}\sum_{h}\hat{a}_{ih}^{2}\right)\rfloor-\lfloor\log_{2}\max_{h}|\hat{a}_{ih}|\rfloor},$
$\nu_{j}$ $\leq 2^{\lfloor\frac{\log_{2}(P-1)-1}{2}-1-\max\left(1,\ 0.5\cdot\log_{2}\sum_{h}\hat{a}_{ih}^{2}\right)\rfloor-\lfloor\log_{2}\max_{h}|\hat{b}_{hj}|\rfloor}.$
For any real number $x\in\mathbb{R}$, $\lfloor\log_{2}|x|\rfloor$ represents the Unit in the First Place (UFP) of $x$, and thus it can be obtained directly using the $\mathtt{ilogb}(\cdot)$ function. Following [uchino_ozaki2], the quantity $(\log_{2}(P-1)-1)/2-1$ is precomputed as $\mathcal{P}^{\prime}_{\mathit{fast}}:=\mathrm{single}(\log_{2}(P-1)/2-1.5)$. Furthermore, $\log_{2}\|\hat{a}_{i,:}\|_{2}$ and $\log_{2}\|\hat{b}_{:,j}\|_{2}$ can be upper-bounded by $0.51\cdot\mathtt{\_\_log2f}(\sum_{h=1}^{2k}\hat{a}_{ih}^{2})$ and $0.51\cdot\mathtt{\_\_log2f}(\sum_{h=1}^{2k}\hat{b}_{hj}^{2})$, respectively, using floating-point arithmetic in round-up mode. Here, $\mathtt{\_\_log2f}$ denotes the fast built-in base-2 logarithm function provided by both the CUDA Math API and the HIP math API. According to NVIDIA’s official CUDA documentation, the maximum absolute error of $\mathtt{\_\_log2f}(x)$ is $2^{-22}$ for $x\in[0.5,2]$, and at most 2 Units in the Last Place (ULP) otherwise [cuda]. The definition of ULP used here follows Muller [muller:inria-00070503]. In addition, the AMD HIP documentation states that the maximum absolute error of $\mathtt{\_\_log2f}(x)$ is 1 ULP for $x\in[10^{-6},10^{6}]$ [hip], although the precise definition of ULP is not specified. Based on these specifications, we assume the bound $\mathtt{\_\_log2f}(x)\leq(\log_{2}x)/(1-4u)$, where $u$ is the unit roundoff of FP32 arithmetic. If $x$ is not sufficiently close to $1$, it is first scaled to $x^{\prime}:=x/2^{\lfloor\log_{2}|x|\rfloor}$, after which the logarithm is evaluated as $\mathtt{\_\_log2f}(x^{\prime})+\lfloor\log_{2}|x|\rfloor$. Based on the above considerations, for $\delta:=\mathrm{single}(0.5/(1-4u))$ computed in round-down mode, the scaling vectors $\mu$ and $\nu$ are determined as follows:
$\mu_{i}$ $:=2^{\lfloor\mathcal{P}^{\prime}_{\mathit{fast}}-\max(1,\ \delta\cdot\mathtt{\_\_log2f}(\sum_{h}\hat{a}_{ih}^{2}))\rfloor-\lfloor\log_{2}\max_{h}|\hat{a}_{ih}|\rfloor},$ (11)
$\nu_{j}$ $:=2^{\lfloor\mathcal{P}^{\prime}_{\mathit{fast}}-\max(1,\ \delta\cdot\mathtt{\_\_log2f}(\sum_{h}\hat{b}_{hj}^{2}))\rfloor-\lfloor\log_{2}\max_{h}|\hat{b}_{hj}|\rfloor},$ (12)
where $\mathcal{P}^{\prime}_{\mathit{fast}}-\delta\cdot x$ for $x\in\mathbb{R}$ is computed using fused multiply-add (fma) operation in round-down mode.
In accurate mode, first, we apply diagonal scaling to $A_{R}$, $A_{I}$, $B_{R}$, and $B_{I}$ so that the integer parts of the maximum absolute values in each row of $A_{R}$ and $A_{I}$, and in each column of $B_{R}$ and $B_{I}$, fit within 6 bits. Specifically, we scale them as $\mathrm{diag}(\bar{\mu})A_{R}$, $\mathrm{diag}(\bar{\mu})A_{I}$, $B_{R}\mathrm{diag}(\bar{\nu})$, $B_{I}\mathrm{diag}(\bar{\nu})$, and then round them upward using the ceiling operation to obtain 7-bit or smaller integer matrices $\bar{A}_{R},\ \bar{A}_{I},\ \bar{B}_{R},\ \bar{B}_{I}$, respectively. These matrices satisfy
$|\bar{\mu}_{i}(A_{R})_{ij}|$ $\leq(\bar{A}_{R})_{ij},$ $\quad|\bar{\mu}_{i}(A_{I})_{ij}|$ $\leq(\bar{A}_{I})_{ij},$
$|(B_{R})_{ij}\bar{\nu}_{j}|$ $\leq(\bar{B}_{R})_{ij},$ $\quad|(B_{I})_{ij}\bar{\nu}_{j}|$ $\leq(\bar{B}_{I})_{ij}.$
After this preprocessing, we compute the upper bound $\bar{C}_{R}+i\bar{C}_{I}$ of the INT8 matrix product $(\bar{A}_{R}+i\bar{A}_{I})(\bar{B}_{R}+i\bar{B}_{I})$ as follows:
$\bar{C}_{I}$ $:=\bar{A}_{I}\bar{B}_{R}+\bar{A}_{R}\bar{B}_{I},$
$\bar{C}_{R}$ $:=\bar{C}_{I}+(\bar{A}_{R}-\bar{A}_{I})(\bar{B}_{R}-\bar{B}_{I})=\bar{A}_{R}\bar{B}_{R}+\bar{A}_{I}\bar{B}_{I},$
where no computing error occurs in $\bar{A}_{R}-\bar{A}_{I}$ and $\bar{B}_{R}-\bar{B}_{I}$ in INT8. Then, the elements of $\mu$ and $\nu$ can be determined in the same manner as in the real-valued Ozaki-II scheme [uchino_ozaki2], i.e.,
$\mu_{i}$ $:=\bar{\mu}_{i}\cdot 2^{\lfloor\mathcal{P}^{\prime}_{\mathit{accu}}-\delta\cdot\log_{2}(\max_{h}\max((\bar{C}_{R})_{ih},(\bar{C}_{I})_{ih}))\rfloor},$ (13)
$\nu_{j}$ $:=\bar{\nu}_{j}\cdot 2^{\lfloor\mathcal{P}^{\prime}_{\mathit{accu}}-\delta\cdot\log_{2}(\max_{h}\max((\bar{C}_{R})_{hj},(\bar{C}_{I})_{hj}))\rfloor},$ (14)
where $\mathcal{P}^{\prime}_{\mathit{accu}}:=\mathrm{single}(\log_{2}(P-1)/2-0.5)$, and $\mathcal{P}^{\prime}_{\mathit{accu}}-\delta\cdot x$ for $x\in\mathbb{R}$ is computed using fma in round-down mode.
III-C Performance Analysis
To analyze the performance characteristics of the proposed method, we employ a simple yet expressive performance model that captures both computation and memory-transfer costs. Let $b$ denote the peak sustained memory bandwidth (in Byte/s) and $p$ denote the peak throughput of the INT8 matrix-multiplication engine (in OPS). For a given operation, the memory-bound execution time is estimated as the required data movement divided by $b$, while the compute-bound time is obtained as the arithmetic workload divided by $p$. Because even memory-bound kernels incur a non-negligible amount of arithmetic overhead, we include a correction term $c$.
III-C1 Fast Mode
Let the input matrices $A\in\mathbb{C}^{m\times k}$ and $B\in\mathbb{C}^{k\times n}$ be double precision. In step 1, both $A$ and $B$ are loaded twice: once for computing the rowwise maximum absolute values of $\hat{A}$ and the columnwise maximum absolute values of $\hat{B}$, and once for computing the residues of $\hat{A}$ and $\hat{B}$. Afterward, the scaling vectors and the INT8 matrices are stored. Each scaling factor is a power of two, so only its exponent needs to be stored as an INT16. Including the correction term $c$, which accounts for arithmetic overhead even in memory-bound operations, the total execution time of step 1 can be estimated as $((32+3N+c)k+2)(m+n)/b$. In step 2, the INT8 GEMM kernels are executed $3N$ times. The resulting $3N$ INT32 matrices are then loaded and converted into $N$ INT8 complex matrices. Accounting for computational overhead $c$, the execution time of step 2 is estimated as $6Nmnk/p+(14N+c)mn/b$. In step 3, the $N$ INT8 complex matrices are loaded, and the CRT accumulation and final reduction are performed. The scaling vectors are then loaded, the inverse scaling is applied, and finally the resulting double-precision complex matrix is stored. With the correction term $c$, the execution time of step 3 is estimated as $(2(m+n)+(2N+16+c)mn)/b$.
Combining the costs derived for steps 1, 2, and 3, the total execution time of the proposed method can be estimated as the sum of its memory-bound and compute-bound components. Aggregating all terms yields the following expression for the overall execution time:
$\begin{multlined}\frac{((3N+32+c)k+4)(m+n)+(16N+16+2c)mn}{b}\\ +\frac{6Nmnk}{p}.\end{multlined}\frac${((3N+32+c)k+4)(m+n)+(16N+16+2c)mn}{b}\\ $\frac${6Nmnk}{p}.
Similarly, when both $A$ and $B$ are single-precision complex matrices, the total execution time can be estimated as
$\begin{multlined}\frac{((3N+16+c)k+4)(m+n)+(16N+8+2c)mn}{b}\\ +\frac{6Nmnk}{p}.\end{multlined}\frac${((3N+16+c)k+4)(m+n)+(16N+8+2c)mn}{b}\\ $\frac${6Nmnk}{p}.
III-C2 Accurate Mode
In the accurate mode, an additional INT8 complex matrix multiplication is performed in step 1, which also increases the memory traffic due to extra stores of INT8 matrices and loads of INT32 matrices, as well as additional loads and stores of the scaling vectors. The total execution time in this mode can therefore be estimated as
$\begin{multlined}\frac{((35+3N+c)k+8)(m+n)+(16N+40+2c)mn}{b}\\ +\frac{6(N+1)mnk}{p},\end{multlined}\frac${((35+3N+c)k+8)(m+n)+(16N+40+2c)mn}{b}\\ $\frac${6(N+1)mnk}{p},
and for the single-precision case,
$\begin{multlined}\frac{((19+3N+c)k+8)(m+n)+(16N+32+2c)mn}{b}\\ +\frac{6(N+1)mnk}{p}.\end{multlined}\frac${((19+3N+c)k+8)(m+n)+(16N+32+2c)mn}{b}\\ $\frac${6(N+1)mnk}{p}.
Figs. 2 and 3 show performance-model heatmaps for the fast and accurate modes in single and double precision for $6$ and $13$ moduli, respectively. The TFLOPS was computed as
$\frac{8mnk}{t}\cdot 10^{-12},$
where $t$ is the execution time in seconds. For example, assuming an effective memory bandwidth of 2–4 TByte/s and an observed INT8 GEMM kernel throughput of 1500 TFLOPS on the GH200, the model predicts that ZGEMM emulation in the accurate mode would achieve approximately 120 TFLOPS.
Figure 2: Performance-model heatmaps for the single-precision complex matrix multiplication emulation. The left panel corresponds to the fast mode and the right panel to the accurate mode. The horizontal axis denotes the attainable memory bandwidth, and the vertical axis denotes the achievable INT8 GEMM throughput. The problem size is fixed to $m=n=k=16384$, and the correction term is set to $c=6$, equal to the number of moduli used in the emulation. The color scale indicates the predicted throughput in TFLOPS of the proposed emulation.
Figure 3: Performance-model heatmaps for the double-precision complex matrix multiplication emulation. The left panel corresponds to the fast mode and the right panel to the accurate mode. The horizontal axis denotes the attainable memory bandwidth, and the vertical axis denotes the achievable INT8 GEMM throughput. The problem size is fixed to $m=n=k=16384$, and the correction term is set to $c=13$, equal to the number of moduli used in the emulation. The color scale indicates the predicted throughput in TFLOPS of the proposed emulation.
III-D GEMM Emulation Library
In this study, we developed a library that provides a complete emulation environment for GEMM operations based on the Ozaki-II scheme. Details of the library and usage instructions are provided in [GEMMul8]. The library supports both CUDA and HIP backends, and the CUDA-based source code is automatically translated into HIP C++ when compiled under the HIP backend. The implementation provides emulation routines for SGEMM and DGEMM as proposed in previous work, as well as the CGEMM and ZGEMM emulation kernels developed in the present study. The supported data types are float, double, cuComplex, cuDoubleComplex, hipFloatComplex, hipDoubleComplex.
To enable transparent integration with existing HPC applications, the library can be activated by adding the path of the shared library to the environment variable LD_PRELOAD. Once activated, the following GEMM calls to CUDA or HIP backends are dynamically redirected to the emulation routines, allowing users to evaluate GEMM emulation without modifying their applications:
  • cublasSgemm, cublasDgemm
  • cublasCgemm, cublasZgemm
  • cublasSgemm_v2, cublasDgemm_v2
  • cublasCgemm_v2, cublasZgemm_v2
  • cublasGemmEx
  • hipblasSgemm, hipblasDgemm
  • hipblasCgemm, hipblasZgemm
  • hipblasGemmEx, hipblasGemmEx_v2
The implementation consists of INT8 matrix multiplications, fixed-order basic arithmetic operations, and the evaluation of the $\log_{2}(\cdot)$ function appearing in (11), (12), (13), and (14). Rounding errors arise only in the computation of the scaling vectors defined by (11), (12), (13), and (14), and in the accumulation and final reduction stages of the CRT. Except for the logarithm evaluation, all computations are deterministic sequences of basic arithmetic, and thus the emulation produces bitwise-reproducible results as long as the precision of the $\log_{2}(\cdot)$ computation is identical. When the API versions coincide, each backend employs identical implementations of __log2f, ensuring reproducibility across runs.
IV Numerical Results
Here, we compare the following methods with each other:
Cgemm:
cublasCgemm or hipblasCgemm using native FP32
Zgemm:
cublasZgemm or hipblasZgemm using native FP64
Zgemm3m:
cublasZgemm3m using native FP64
fast-NN:
Ozaki-II in fast mode with $N$ moduli
accu-NN:
Ozaki-II in accurate mode with $N$ moduli
OS I-$S$:
Ozaki-I from cuBLAS in Fixed Mantissa Control mode with $S$ slices
The hipBLAS does not provide hipblasCgemm3m or hipblasZgemm3m and the same limitation applies to the rocBLAS, though it is not used in this study. We have empirically confirmed that both the performance and numerical accuracy of cublasCgemm3m are essentially identical to those of cublasCgemm; thus, we omit its results from the following comparisons. The cuBLAS implementation of Ozaki-I-based emulation provides two execution modes: Dynamic Mantissa Control and Fixed Mantissa Control. In the Dynamic Mantissa Control mode, the routine determines the number of mantissa bits, equivalently, the number of slices, required for fixed-point emulation such that the numerical accuracy matches or exceeds that of FP64. If the required mantissa length exceeds either an internal limit or a user-specified maximum, the computation is automatically executed in native FP64. In the Fixed Mantissa Control mode, the user specifies the mantissa-bit length used for the fixed-point representation. This mode reduces runtime overhead by eliminating dynamic precision selection, but it does not ensure FP64-level accuracy. In this study, the Fixed Mantissa Control mode is used to enable a direct comparison of computational throughput and to quantify the performance gain attributable to fixed-precision emulation.
IV-A Accuracy
Each element of the test matrices $A\in\mathbb{C}^{m\times k}$ and $B\in\mathbb{C}^{k\times n}$ was generated by independently assigning its real and imaginary parts as
$(\mathrm{rand}-0.5)\cdot\exp(\mathrm{randn}\cdot\phi),$
where $\mathrm{rand}\in(0,1]$ denotes a uniform random number and $\mathrm{randn}$ denotes a standard normal random number. Both types of random numbers were generated using either the cuRAND API or the hipRAND library. Following the methodology adopted in prior studies about the Ozaki schemes [ozaki2012error, ozaki2013generalization, mukunoki2020, ootomo2024dgemm, uchino2025Performance, ozaki-scheme2, uchino_ozaki2], we introduce the parameter $\phi$ to control the dynamic range of the input matrices. In accordance with these studies, we specify $\phi\in\{0,0.5,1,1.5\}$ for single-precision tests and $\phi\in\{0.5,1,2,4\}$ for double-precision tests. As $\phi$ increases, the maximum magnitude grows exponentially, whereas the median and interquartile range remain comparatively stable. This confirms that $\phi$ effectively widens the dynamic range without substantially affecting the central tendency of the distribution. Such controlled variation is essential for evaluating the performance and numerical stability of the Ozaki schemes.
Figs. 4 and 5 show the accuracy of CGEMM and ZGEMM, respectively, observed on the GH200. Similar results were obtained on other GPUs. The maximum relative error in the figures were calculated as
$\max_{i,j}\left(\max\left(\frac{|\mathrm{Re}(\tilde{c}_{ij})-\mathrm{Re}(c_{ij})|}{|\mathrm{Re}(c_{ij})|},\frac{|\mathrm{Im}(\tilde{c}_{ij})-\mathrm{Im}(c_{ij})|}{|\mathrm{Im}(c_{ij})|}\right)\right),$
where $\tilde{C}\in\mathbb{C}^{m\times n}$ is the approximation of $AB$ obtained using the emulation methods and the standard complex GEMM routines, and $C\in\mathbb{C}^{m\times n}$ is the high-precision result of $AB$ computed using double-double arithmetic.
In CGEMM emulation, fast-$N$ with $6\leq N\leq 9$ and accu-$N$ with $6\leq N\leq 8$ provided the accuracy comparable with that of CGEMM. In ZGEMM emulation, fast-$N$ with $13\leq N\leq 18$ and accu-$N$ with $13\leq N\leq 17$ achieved the accuracy comparable with that of ZGEMM. As $\phi$ increases, the range of exponents for $A$ and $B$ expands, making it easier to overestimate the upper bound $2\sum_{h=1}^{k}|a^{\prime}_{ih}||b^{\prime}_{hj}|$ in the condition (4). DGEMM emulation in [uchino_ozaki2] required at least 14 moduli to achieve DGEMM-level accuracy, whereas the proposed method attains ZGEMM-level accuracy with only 13 moduli. This reduction is attributed to the use of the Karatsuba algorithm in the INT8 complex matrix multiplication, which suppresses precision loss more effectively than floating-point arithmetic.
The final reduction $\mathrm{mod}(s_{ij},P)$ of the CRT is defined as $s_{ij}-P\cdot\mathrm{round}(s_{ij}/P)$. This formulation entails subtracting nearly identical quantities, leading to loss of significance. Therefore, obtaining the target accuracy requires that both the modulo operation and the calculation of $s_{ij}$ be conducted at a precision level exceeding that of the final output. Regarding implementation, the limiting accuracy is currently constrained by the arithmetic precision employed: CGEMM emulation relies on double-precision arithmetic, while ZGEMM emulation calculates $s_{ij}$ using the high-precision approach in (5) and employs a simplified high-precision method based on double-double arithmetic for the modulo operation. Enhancing the limiting accuracy therefore necessitates performing these operations with even higher precision.
Figure 4: Maximum relative error of single-precision complex matrix multiplication on NVIDIA GH200 Grace Hopper Superchip, CUDA Toolkit 13.0.88, and gcc 11.5.0, for $m=n=1024$ and $k=16384$.
Figure 5: Maximum relative error of double-precision complex matrix multiplication on NVIDIA GH200 Grace Hopper Superchip, CUDA Toolkit 13.0.88, and gcc 11.5.0, for $m=n=1024$ and $k=16384$.
IV-B Performance
[…]
Figure 6: Throughput performance in TFLOPS of single-precision complex matrix multiplication on NVIDIA GH200 Grace Hopper Superchip, CUDA Toolkit 13.0.88, and gcc 11.5.0, for $m=k=n$. Methods shown in red italic achieve CGEMM-level accuracy.
Figure 7: Throughput performance in TFLOPS of single-precision complex matrix multiplication on NVIDIA B200 Blackwell GPU with Intel Xeon Platinum 8570 CPU, CUDA Toolkit 12.8.93, and g++ 13.3.0, for $m=k=n$. Methods shown in red italic achieve CGEMM-level accuracy.
Figure 8: Throughput performance in TFLOPS of single-precision complex matrix multiplication on NVIDIA GeForce RTX 5080 GPU with AMD Ryzen 7 3700X CPU, CUDA Toolkit 13.0.88, and gcc 11.5.0, for $m=k=n$. Methods shown in red italic achieve CGEMM-level accuracy.
Figure 9: Throughput performance in TFLOPS of single-precision complex matrix multiplication on AMD Instinct MI300X GPU with AMD EPYC 9534 CPU, ROCm 6.4.2, and clang 19.0.0git, for $m=k=n$. Methods shown in red italic achieve CGEMM-level accuracy.
Figure 10: Throughput performance in TFLOPS of double-precision complex matrix multiplication on NVIDIA GH200 Grace Hopper Superchip, CUDA Toolkit 13.0.88, and gcc 11.5.0, for $m=k=n$. Methods shown in red italic achieve ZGEMM-level accuracy.
Figure 11: Throughput performance in TFLOPS of double-precision complex matrix multiplication on NVIDIA B200 Blackwell GPU with Intel Xeon Platinum 8570 CPU, CUDA Toolkit 12.8.93, and g++ 13.3.0, for $m=k=n$. Methods shown in red italic achieve ZGEMM-level accuracy.
Figure 12: Throughput performance in TFLOPS of double-precision complex matrix multiplication on NVIDIA GeForce RTX 5080 GPU with AMD Ryzen 7 3700X CPU, CUDA Toolkit 13.0.88, and gcc 11.5.0, for $m=k=n$. Methods shown in red italic achieve ZGEMM-level accuracy.
Figure 13: Throughput performance in TFLOPS of double-precision complex matrix multiplication on AMD Instinct MI300X GPU with AMD EPYC 9534 CPU, ROCm 6.4.2, and clang 19.0.0git, for $m=k=n$. Methods shown in red italic achieve ZGEMM-level accuracy.
These performance characteristics correlate strongly with the architectural balance of arithmetic throughput on each GPU. The GH200 provides a ratio of $\text{FP64}:\text{FP32}:\text{INT8}=1:1:29.5$, indicating that INT8 operations are moderately more abundant than FP64 and FP32 units. The B200 exhibits an even more pronounced imbalance, $1:2:121.6$, where INT8 throughput overwhelmingly dominates the floating-point pipelines. This architectural skew directly amplifies the benefit of the proposed kernels, which replace large portions of floating-point arithmetic with INT8-intensive computation. In contrast, the MI300X offers a more subdued ratio of $1:1:16$, providing substantially less INT8 advantage relative to its floating-point throughput. As a result, the computational substitution enabled by our method yields only marginal acceleration on the MI300X, consistent with the limited speedups observed in the benchmarks. For the RTX 5080, the observed performance gains, especially the very large speedups in ZGEMM, suggest a much stronger favoring of INT8 throughput relative to floating-point units. The architectural balance on RTX 5080 corresponds to $\text{FP64}:\text{FP32}:\text{INT8}=1:64:512$, which explains the pronounced acceleration when using INT8-based emulation.
Compared with the Ozaki-I emulation implemented in cuBLAS, the proposed method delivered substantially higher throughput on both GH200 and RTX 5080 GPUs. On the GH200, Ozaki-I with 7–9 slices achieved 25–39 TFLOPS, whereas the proposed method attained 99–136 TFLOPS. On the RTX 5080, Ozaki-I achieved 6.4–10 TFLOPS, while the proposed method reached 19–26 TFLOPS. These disparities arise from the fundamental algorithmic differences: Ozaki-I with $S$ slices requires $S(S+1)/2$ INT8 complex matrix multiplications, whereas the proposed technique with $N$ moduli requires only $N$ INT8 complex matrix multiplications and and further reduces their cost by a factor of $3/4$ using the Karatsuba method. This reduction in both the number and cost of INT8 matrix multiplication kernel invocations directly explains the substantially higher performance of the proposed method.
IV-C Supplemental Comparison for Real-Valued Emulation
Here, we focus on double-precision real matrix multiplication. In this section, we focus on double-precision real matrix multiplication. We compare two emulation strategies: the Ozaki-I based emulation provided by cuBLAS and the Ozaki-II based emulation proposed in [uchino_ozaki2], enhanced with the blocking technique introduced in Section III. This supplemental analysis aims to isolate the behavior of Ozaki-type schemes in the real domain and to clarify the performance implications of incorporating blocking into Ozaki-II under conditions consistent with real-valued DGEMM emulation.
According to [uchino_ozaki2], the fast mode and the accurate mode require 14–18 and 14–17 moduli, respectively, to achieve DGEMM-level accuracy. On the GH200 with matrix size $m=n=k=16384$, the Ozaki-II fast-$N$ configuration with $14\leq N\leq 18$ achieved 63–82 TFLOPS, while the accu-$N$ configuration with $14\leq N\leq 17$ reached 63–75 TFLOPS. Incorporating the blocking technique improved performance across all settings: the blocked fast-$N$ version attained 72–93 TFLOPS, and the blocked accu-$N$ version delivered 71–85 TFLOPS. cuBLAS’s Ozaki-I implementation with 7–10 slices achieved only 20–39 TFLOPS, and the standard DGEMM kernel sustained 61 TFLOPS.
On the RTX 5080 with $m=n=k=8192$, the fast-$N$ and accu-$N$ configurations of the proposed method reached 12–16 TFLOPS and 12–15 TFLOPS, respectively, while cuBLAS’s Ozaki-I achieved 5.4–10 TFLOPS. In comparison, the native DGEMM kernel delivered 0.84 TFLOPS, highlighting the substantial advantage of INT8-intensive Ozaki-type emulation on this consumer architecture.
These results demonstrate that blocking substantially enhances the performance of Ozaki-II for large-scale problems, particularly on bandwidth-rich GPUs such as the H200, and Ozaki-II, whether blocked or unblocked, consistently outperforms cuBLAS’s Ozaki-I due to its reduced number of INT8 matrix multiplication invocations.
V Conclusion
This study proposed emulation methods of complex matrix multiplication based on the Ozaki-II scheme and developed a portable library that supports execution on both AMD and NVIDIA GPUs. The numerical results demonstrated that the proposed methods outperform the standard GEMM routines in cuBLAS and hipBLAS, which are implemented using native FP32 and FP64 arithmetic. In addition, the proposed methods outperform the emulation using the Ozaki-I scheme implemented in cuBLAS.
Future work includes comparisons with the BF16x9 CGEMM emulation algorithm provided for NVIDIA B200 and B300 GPUs in cuBLAS 12.9 and later. A major limitation of the proposed method, as well as emulation-based approaches in general, is the substantial working memory required. This overhead is currently unavoidable when aiming to achieve high performance without relying on FP32 or FP64 matrix multiplication. Addressing this issue remains an open challenge not only for emulation techniques but also for HPC applications that employ them.
The Ozaki-I and Ozaki-II schemes utilize INT8 matrix engines, benefiting from the fact that the error-free low-precision matrix multiplications appearing in these schemes do not require exponent handling, enabling efficient synergy with INT8 operations with INT32 accumulation. When using FP16 or BF16 matrix engines with FP32 accumulation, the effective representable integer range is limited to 24 bits due to the FP32 significand, leading to significantly reduced computational efficiency compared with INT32 accumulation. However, some recent GPUs, such as the NVIDIA B300, provide reduced INT8 capabilities. Supporting such hardware architectures represents an additional direction for future work.
Acknowledgment
We thank Prof. Rio Yokota for his cooperation in conducting the experiments on the B200. This study was supported by Japan Society for the Promotion of Science Grant-in-Aid Numbers 25K03126 (Scientific Research B) and 24K23874 (Research Activity Start-up).

### v3

@@
 General-purpose computing architectures have increasingly adopted low-precision matrix multiplication units in graphics processing units (GPUs), tensor processing units (TPUs), and neural processing units (NPUs).
-Representative examples include Tensor Cores in NVIDIA GPUs [tensorcore], Matrix Cores in AMD GPUs [matrixcore], the MXU in Google TPUs [tpu], and the MAC array in Intel NPUs [intelXecore].
-These units offer substantially higher arithmetic throughput than high-precision units and have become central components of modern accelerator design.
-Table I shows the theoretical throughput performance in TFLOPS or TOPS of AMD [MI300A, MI300X, MI325X, MI350X, MI355X] and NVIDIA GPUs [A100, H100, Blackwell, RTX5080].
-In recent years, such units have also begun to appear as standard features even in laptop-class processors [copilot+pc, SME], further broadening their availability.
-While these units were originally motivated by machine-learning workloads, their architectural characteristics also present an opportunity to accelerate high-precision computations by emulating them on low-precision hardware.
+Representative examples include Tensor Cores in NVIDIA GPUs [tensorcore], Matrix Cores in AMD GPUs [matrixcore], Matrix Multiply Units in Google TPUs [tpu], and Multiply-Accumulate arrays in Intel NPUs [intelXecore].
+These units offer substantially higher arithmetic throughput than that of high-precision units and have become central components of modern accelerator design.
+Table I shows the theoretical throughput performance for AMD [MI300A, MI300X, MI325X, MI350X, MI355X] and NVIDIA GPUs [A100, H100, Blackwell, RTX5080].
+In recent years, low-precision matrix multiplication units have also begun to appear as standard features even in laptop-class processors [copilot+pc, SME], further broadening their availability.
+While these units were originally motivated by machine-learning workloads, their architectural characteristics also present an opportunity to accelerate high-precision computations by emulating such computations on low-precision hardware.
 Consequently, emulation-based high-precision matrix multiplication has emerged as an active research topic in the high-performance computing community [brower2025mixedprecisionabinitiotensor, abdelfattah2025analysisfloatingpointmatrixmultiplication, schwarz2025guaranteed, mukunoki2025sparseiterativesolversusing, dongarra2024hardwaretrendsimpactingfloatingpoint, 11196192, 11196413].
-TABLE I: TFLOPS and TOPS of AMD and NVIDIA GPUs [MI300A, MI300X, MI325X, MI350X, MI355X, A100, H100, Blackwell, RTX5080]
+TABLE I: Performance (in TFLOPS or TOPS where appropriate) of AMD and NVIDIA GPUs [MI300A, MI300X, MI325X, MI350X, MI355X, A100, H100, Blackwell, RTX5080]
 GPU FP64 FP32 FP16 FP8 INT8
@@
 A prior line of work introduced the Ozaki-I [ozaki2012error, ozaki2013generalization, mukunoki2020, mukunoki2025dgemmfp64arithmetic, ootomo2024dgemm, uchino2025Performance] and Ozaki-II [ozaki-scheme2, uchino_ozaki2] schemes as frameworks for emulating high-precision matrix multiplication using low-precision matrix engines, demonstrating high performance and power efficiency on contemporary GPUs.
-In parallel, the Ozaki-I scheme has recently been incorporated into cuBLAS (from version 13.0 Update 2), and this implementation supports complex matrix multiplication.
+The Ozaki-I scheme has recently been incorporated into cuBLAS (from version 13.0 Update 2).
+This implementation supports complex matrix multiplication.
 However, for real-valued matrix multiplication, previous studies have reported that the Ozaki-II scheme achieves higher performance than Ozaki-I.
 This suggests that a complex extension of the Ozaki-II scheme would likewise outperform the complex Ozaki-I approach.
-Despite this expectation, no complex-valued formulation of Ozaki-II has been developed to date.
-This absence is notable because complex matrix multiplication is a key computational kernel in many HPC applications, and its performance can dominate the overall runtime of large-scale scientific simulations.
+Despite this expectation, no complex-valued formulation of the Ozaki-II scheme has been developed.
+This absence is notable because complex matrix multiplication is a key computational kernel in many high-performance computing applications and its performance can dominate the overall runtime of large-scale scientific simulations.
 To address this gap, the present work extends the Ozaki-II scheme to single- and double-precision complex matrix multiplication (CGEMM and ZGEMM, respectively).
 We develop emulation algorithms for CGEMM and ZGEMM by combining the Ozaki-II scheme with a Karatsuba-based formulation of complex arithmetic.
-The resulting methods maintain numerical accuracy comparable to standard CGEMM and ZGEMM while effectively exploiting the high throughput of low-precision matrix engines.
-In addition, several improvements applicable to existing SGEMM and DGEMM emulation techniques are introduced, further enhancing the efficiency of the Ozaki schemes in the real domain.
-The proposed algorithms are released as a portable library supporting both CUDA and HIP backends, enabling execution on NVIDIA GPUs and AMD GPUs.
-Performance is evaluated on an NVIDIA GH200 Grace Hopper Superchip, an NVIDIA B200 Blackwell GPU, an NVIDIA GeForce RTX 5080 Blackwell GPU, and an AMD Instinct MI300X GPU, and compared against cuBLAS and hipBLAS implementations.
-For sufficiently large matrix sizes, the proposed methods surpass both vendor-optimized libraries and, in particular, achieve higher throughput than the cuBLAS Ozaki-I–based ZGEMM emulation, demonstrating the performance potential of Ozaki-II-based emulation for CGEMM and ZGEMM.
-We also provide performance models for the proposed methods, and the predicted throughput closely aligns with the measured results across all tested configurations.
+The developed methods maintain numerical accuracy comparable to that of standard CGEMM and ZGEMM while effectively exploiting the high throughput of low-precision matrix engines.
+In addition, several improvements applicable to existing single- and double-precision general matrix multiplication (GEMM) emulation techniques are introduced, further enhancing the efficiency of the Ozaki schemes in the real domain.
+The proposed algorithms are released as a portable library that supports both CUDA and HIP backends, enabling execution on NVIDIA and AMD GPUs.
+Performance is evaluated on an NVIDIA GH200 Grace Hopper Superchip, an NVIDIA B200 Blackwell GPU, an NVIDIA GeForce RTX 5080 Blackwell GPU, and an AMD Instinct MI300X GPU and compared against cuBLAS and hipBLAS implementations.
+For sufficiently large matrix sizes, the proposed methods surpass both vendor-optimized libraries and, in particular, achieve higher throughput than that of cuBLAS Ozaki-I-based ZGEMM emulation, demonstrating the performance potential of Ozaki-II-based emulation for CGEMM and ZGEMM.
+We also provide performance models for the proposed methods.
+The predicted throughput closely aligns with the measurement results across all tested configurations.
 The remainder of this paper is organized as follows.
@@
 Section III presents the proposed efficient methods for emulating CGEMM and ZGEMM.
-Section IV shows the numerical results comparing the proposed emulation with the vender-optimized GEMM routines.
+Section IV shows numerical results comparing the proposed emulation with vender-optimized GEMM routines.
 Section V presents the concluding remarks.
-Note that in this paper, we do not consider matrix shapes for which performance becomes memory-bound, such as small matrices or highly skinny matrices, because low-precision matrix engines do not achieve high throughput on such workloads.
+Note that in this paper, we do not consider matrix shapes for which performance becomes memory-bound, such as small matrices or very skinny matrices, because low-precision matrix engines do not achieve high throughput on such workloads.
 II Ozaki-II Scheme for Real Matrix Multiplication
-The Ozaki-II scheme is an emulation method of matrix multiplication based on a large-integer multiplication technique using the Chinese Remainder Theorem (CRT).
-The scheme for matrix multiplication of $A\in\mathbb{R}^{m\times k}$ and $B\in\mathbb{R}^{k\times n}$ consists of the following three steps:
+The Ozaki-II scheme is an emulation method for matrix multiplication based on a large-integer multiplication technique using the Chinese remainder theorem (CRT).
+The scheme for the multiplication of matrices $A\in\mathbb{R}^{m\times k}$ and $B\in\mathbb{R}^{k\times n}$ consists of the following three steps:
   1. 1.
@@
   3. 3.
-    Convert $C^{\prime}$ back to floating-point format $C\in\mathbb{R}^{m\times n}$ by inversely scaling.
+    Convert $C^{\prime}$ back to floating-point format $C\in\mathbb{R}^{m\times n}$ via inverse scaling.
 The integer matrix multiplication in step 2 is performed by decomposing the elements of $A^{\prime}$ and $B^{\prime}$ into residues with respect to a predefined set of pairwise coprime moduli.
 Let $p_{1},p_{2},\dots,p_{N}\in\mathbb{N}$ denote these moduli and define $P:=\prod_{\ell=1}^{N}p_{\ell}$.
-Let $\mathrm{mod}(x,p_{\ell})$ for $x\in\mathbb{Z}$ denote the symmetric remainder of $x/p_{\ell}$, i.e., $\mathrm{mod}(x,p_{\ell})=x-p_{\ell}\cdot\mathrm{round}(x/p_{\ell})$, where $\mathrm{round}(\cdot)$ rounds its input to the nearest integer.
+Let $\mathrm{mod}(x,p_{\ell})$ for $x\in\mathbb{Z}$ denote the symmetric remainder of $x/p_{\ell}$ (i.e., $\mathrm{mod}(x,p_{\ell})=x-p_{\ell}\cdot\mathrm{round}(x/p_{\ell})$, where $\mathrm{round}(\cdot)$ rounds its input to the nearest integer).
 Each element of $A^{\prime}$ and $B^{\prime}$ is mapped to its residue representation modulo each $p_{\ell}$ as
@@
 $2\sum_{h=1}^{k}|a^{\prime}_{ih}||b^{\prime}_{hj}|<P$ (4)
-is satisfied for all $(i,j)$ pairs, the possible value of $c^{\prime}_{ij}$ is only $\mathrm{mod}\left((A^{\prime}B^{\prime})_{ij},P\right)$ and the CRT guarantees that the reconstructed $C^{\prime}$ is unique.
+is satisfied for all $(i,j)$ pairs, the only possible value of $c^{\prime}_{ij}$ is $\mathrm{mod}\left((A^{\prime}B^{\prime})_{ij},P\right)$ and the CRT guarantees that the reconstructed $C^{\prime}$ is unique.
 Therefore, the scaling vector used in step 1 must be chosen such that the bound (4) is ensured for $A^{\prime}$ and $B^{\prime}$.
@@
 Hereafter, we assume that $k\leq 2^{17}$ to perform INT8 matrix multiplication without error.
-In [uchino_ozaki2], step iv at the line 7 of Algorithm 1 uses the standard (non-symmetric) modulo operation that returns values in $0,\,p_{\ell})$ and $E^{\prime}_{\ell}$ are UINT8 matrices.
-In DGEMM emulation, step v at the line 7 evaluates $S=\sum_{\ell=1}^{N}P/p_{\ell}\cdot q_{\ell}\cdot E^{\prime}_{\ell}$ by the unevaluated sum $S_{1}+S_{2}$ as
+In [uchino_ozaki2], step V-iv of Algorithm 1 uses the standard (non-symmetric) modulo operation that returns values in $0,\,p_{\ell})$ and $E^{\prime}_{\ell}$ are UINT8 matrices.
+In DGEMM emulation, step V-v of Algorithm 1 evaluates $S=\sum_{\ell=1}^{N}P/p_{\ell}\cdot q_{\ell}\cdot E^{\prime}_{\ell}$ using the unevaluated sum $S_{1}+S_{2}$ as
 $S\approx S_{1}+S_{2},\quad S_{1}:=\sum_{\ell=1}^{N}s_{\ell 1}E^{\prime}_{\ell},\quad S_{2}:=\sum_{\ell=1}^{N}s_{\ell 2}E^{\prime}_{\ell},$ (5)
@@
 This bit allocation ensures that all products $s_{\ell 1}E^{\prime}_{\ell}$ can be accumulated in $S_{1}$ without rounding error.
-In contrast, the present work adopts a symmetric modulo operation in step iv at the line 7, ensuring that each $E^{\prime}_{\ell}$ becomes INT8 from UINT8.
-Consequently, the upper part $s_{\ell 1}$ of $P/p_{\ell}\cdot q_{\ell}$ can use $(53-7-\lceil\log_{2}N\rceil)$ bits instead of $(53-8-\lceil\log_{2}N\rceil)$ bits, gaining one additional effective bit compared with the previous method.
+In contrast, the present work adopts a symmetric modulo operation in step V-iv of Algorithm 1, ensuring that each $E^{\prime}_{\ell}$ becomes INT8 from UINT8.
+Consequently, the upper part $s_{\ell 1}$ of $P/p_{\ell}\cdot q_{\ell}$ can use $(53-7-\lceil\log_{2}N\rceil)$ bits instead of $(53-8-\lceil\log_{2}N\rceil)$ bits, gaining one additional effective bit compared to the previous method.
 This slight but meaningful increase in representable precision improves the numerical robustness of the reconstruction step without changing the overall algorithmic structure.
 Algorithm 1 Outline of Ozaki-II scheme using INT8 matrix engines for real matrix multiplication [uchino_ozaki2]. The function $\mathrm{trunc}(\cdot)$ returns the integer part of its input.
-[…]
-[…]
-3:Determine the moduli $p_{\ell}\leq 256$ for $\ell=1,\dots,N$.
-4:Determine modular multiplicative inverses $q_{\ell}$ of $P/p_{\ell}$ modulo $p_{\ell}$ for $\ell=1,\dots,N$.
-5:Determine the scaling vectors $\mu\in\mathbb{R}^{m}$ and $\nu\in\mathbb{R}^{n}$.
-6:Convert AA and $B$ to $A^{\prime}\in\mathbb{Z}^{m\times k}$ and $B^{\prime}\in\mathbb{Z}^{k\times n}$:
+$A\in\mathbb{R}^{m\times k}$, $B\in\mathbb{R}^{k\times n}$, $N\in\mathbb{N}$
+$C\approx AB\in\mathbb{R}^{m\times n}$
+Step I
+Determine the moduli $p_{\ell}\leq 256$ for $\ell=1,\dots,N$.
+Step II
+Determine modular multiplicative inverses $q_{\ell}$ of $P/p_{\ell}$ modulo $p_{\ell}$ for $\ell=1,\dots,N$.
+Step III
+Determine the scaling vectors $\mu\in\mathbb{R}^{m}$ and $\nu\in\mathbb{R}^{n}$.
+Step IV
+Convert $A$ and $B$ to $A^{\prime}\in\mathbb{Z}^{m\times k}$ and $B^{\prime}\in\mathbb{Z}^{k\times n}$:
   1. i)
@@
     $b^{\prime}_{ij}:=\mathrm{trunc}(b_{ij}\cdot\nu_{j})$
-7: Compute $C^{\prime}\approx A^{\prime}B^{\prime}$ based on the CRT:
+Step V
+Compute $C^{\prime}\approx A^{\prime}B^{\prime}$ based on the CRT:
   1. i)
@@
     $c^{\prime}_{ij}:=\bmod(s_{ij},P)$
-8:Convert $C^{\prime}$ back to floating-point format $C\in\mathbb{R}^{m\times n}$:
+Step VI
+Convert $C^{\prime}$ back to floating-point format $C\in\mathbb{R}^{m\times n}$:
   1. i)
     $C:=\mathrm{diag}(\mu)^{-1}C^{\prime}\mathrm{diag}(\nu)^{-1}$
-To satisfy the condition (4) efficiently, the scaling vectors used in step 1 must be determined appropriately.
+To satisfy the condition (4) efficiently, the scaling vectors used in step IV of Algorithm 1 must be determined appropriately.
 Previous work introduced two strategies for computing this scaling: a fast mode and an accurate mode.
-The fast mode estimates an upper bound of $\sum_{h=1}^{k}|a^{\prime}_{ih}||b^{\prime}_{hj}|$ via the Cauchy–Schwarz inequality and determines the scaling vector analytically based on this estimate.
-This approach incurs minimal preprocessing cost, but the bound obtained from the Cauchy–Schwarz inequality tends to be overly conservative.
-As a consequence, a larger $P$ (i.e., more moduli) is often required to satisfy the uniqueness condition, increasing the computational workload in step 2.
+The fast mode estimates the upper bound of $\sum_{h=1}^{k}|a^{\prime}_{ih}||b^{\prime}_{hj}|$ via the Cauchy–Schwarz inequality and determines the scaling vector analytically based on this estimate.
+This approach incurs minimal preprocessing cost, but the obtained bound tends to be overly conservative.
+As a consequence, a larger $P$ (i.e., more moduli) is often required to satisfy the uniqueness condition, increasing the computational workload in step V Algorithm 1.
 In contrast, the accurate mode estimates the upper bound by performing an auxiliary matrix multiplication using INT8 matrix engines.
-Although this mode requires additional computation during step 1, the direct estimation yields a much tighter bound than that obtained via the Cauchy–Schwarz inequality.
+Although this mode requires additional computation during step III of Algorithm 1, the direct estimation yields a much tighter bound than that obtained via the Cauchy–Schwarz inequality.
 This typically enables the use of fewer moduli, reducing the number of modular matrix multiplications and potentially improving overall performance.
 III Proposed Methods
-This section presents the extension of the Ozaki-II scheme to complex matrix multiplication, targeting high-performance execution on INT8 matrix engines.
-Let $A=A_{R}+iA_{I}\in\mathbb{C}^{m\times k}$ for $A_{R},A_{I}\in\mathbb{R}^{m\times k}$ and $B=B_{R}+iB_{I}\in\mathbb{C}^{k\times n}$ for $B_{R},B_{I}\in\mathbb{R}^{k\times n}$, and we compute $C=C_{R}+iC_{I}:=AB$, where $C_{R},C_{I}\in\mathbb{R}^{m\times n}$.
-A standard approach to handle complex matrix multiplication is to employ the following representation:
+This section presents the proposed extension of the Ozaki-II scheme to complex matrix multiplication, targeting high-performance execution on INT8 matrix engines.
+Let $A=A_{R}+iA_{I}\in\mathbb{C}^{m\times k}$ for $A_{R},A_{I}\in\mathbb{R}^{m\times k}$ and $B=B_{R}+iB_{I}\in\mathbb{C}^{k\times n}$ for $B_{R},B_{I}\in\mathbb{R}^{k\times n}$.
+We compute $C=C_{R}+iC_{I}:=AB$, where $C_{R},C_{I}\in\mathbb{R}^{m\times n}$.
+A standard approach for handling complex matrix multiplication is to employ the following representation:
 $\hat{M}:=\begin{pmatrix}M_{R}&-M_{I}\\ M_{I}&M_{R}\end{pmatrix}$ (6)
@@
 III-A INT8 Complex Matrix Multiplication
-Using (6), $C_{R}$ and $C_{I}$ are obtained by
+Using (6), $C_{R}$ and $C_{I}$ are obtained using
 $\begin{pmatrix}C_{R}\\ C_{I}\end{pmatrix}=\begin{pmatrix}A_{R}&-A_{I}\\ A_{I}&A_{R}\end{pmatrix}\begin{pmatrix}B_{R}\\ B_{I}\end{pmatrix}$ (7)
@@
 For the formulations, the applicability condition on the inner dimension changes from $k\leq 2^{17}$ to $k\leq 2^{16}$.
-Another standard approach to handle complex matrix multiplication is the Karatsuba algorithm.
+Another standard approach for handling complex matrix multiplication is the Karatsuba algorithm.
 The Karatsuba algorithm expresses the complex product using three real matrix multiplications as
@@
 […]
-Figure 1: Performance comparison of four INT8-based GEMM strategies on NVIDIA H100 NVL with CUDA Toolkit 12.8.61.
+Figure 1: Performance comparison of four INT8-based GEMM strategies on NVIDIA H100 NVL (CUDA Toolkit 12.8.61).
 The first two methods use a single INT8 matrix multiplication with dimensions $(2h,h,2h)$ and $(h,2h,2h)$, respectively.
-The third method performs three INT8 multiplications of size $(h,h,h)$, while the fourth applies the same three-multiplication scheme but introduces blocking along the $n$ dimension.
+The third method performs three INT8 multiplications of size $(h,h,h)$.
+The fourth method applies the same three-multiplication scheme but introduces blocking along the $n$ dimension.
 III-B Scaling Vector Determination
-First, we convert $\hat{A}$ and $\hat{B}$ to integer matrices $\hat{A}^{\prime}$ and $\hat{B}^{\prime}$ as
+First, we convert $\hat{A}$ and $\hat{B}$ to integer matrices $\hat{A}^{\prime}$ and $\hat{B}^{\prime}$, respectively, as
 $\hat{a}^{\prime}_{ij}:=\mathrm{trunc}(\hat{a}_{ij}\cdot\mu_{i}),\quad\hat{b}^{\prime}_{ij}:=\mathrm{trunc}(\hat{b}_{ij}\cdot\nu_{j}).$
-We also consider two strategies: a fast mode and an accurate mode. From $|\hat{a}^{\prime}_{ij}|\leq|\hat{a}_{ij}|\cdot\mu_{i}$ and $|\hat{b}^{\prime}_{ij}|\leq|\hat{b}_{ij}|\cdot\nu_{j}$,
+We consider two strategies: a fast mode and an accurate mode. From $|\hat{a}^{\prime}_{ij}|\leq|\hat{a}_{ij}|\cdot\mu_{i}$ and $|\hat{b}^{\prime}_{ij}|\leq|\hat{b}_{ij}|\cdot\nu_{j}$,
 $\sum_{h=1}^{k}|\hat{a}^{\prime}_{ih}||\hat{b}^{\prime}_{hj}|\leq\mu_{i}\nu_{j}\cdot\sum_{h=1}^{k}|\hat{a}_{ih}||\hat{b}_{hj}|.$
@@
 From (6), the row norms of $\hat{A}$ satisfy $\|\hat{a}_{i,:}\|_{2}=\|\hat{a}_{i+m,:}\|_{2}$ for $1\leq i\leq m$, and the column norms of $\hat{B}$ satisfy $\|\hat{b}_{:,j}\|_{2}=\|\hat{b}_{:,j+n}\|_{2}$ for $1\leq j\leq n$.
-Thus, the scaling values of rows $i$ and $i+m$ of $\hat{A}$ are identical, and the same holds for columns $j$ and $j+n$ of $\hat{B}$.
+Thus, the scaling values for rows $i$ and $i+m$ of $\hat{A}$ are identical; the same holds for columns $j$ and $j+n$ of $\hat{B}$.
 Consequently, the scaling vectors reduce to $\mu\in\mathbb{R}^{m}$ and $\nu\in\mathbb{R}^{n}$.
-The elements of $\mu$ and $\nu$ can then be determined in the same manner as in the real-valued Ozaki-II scheme [uchino_ozaki2], i.e.,
+The elements of $\mu$ and $\nu$ can then be determined in the same manner as in the real-valued Ozaki-II scheme [uchino_ozaki2]; that is,
 $\mu_{i}$ $\leq 2^{\lfloor\frac{\log_{2}(P-1)-1}{2}-1-\max\left(1,\ 0.5\cdot\log_{2}\sum_{h}\hat{a}_{ih}^{2}\right)\rfloor-\lfloor\log_{2}\max_{h}|\hat{a}_{ih}|\rfloor},$
 $\nu_{j}$ $\leq 2^{\lfloor\frac{\log_{2}(P-1)-1}{2}-1-\max\left(1,\ 0.5\cdot\log_{2}\sum_{h}\hat{a}_{ih}^{2}\right)\rfloor-\lfloor\log_{2}\max_{h}|\hat{b}_{hj}|\rfloor}.$
-For any real number $x\in\mathbb{R}$, $\lfloor\log_{2}|x|\rfloor$ represents the Unit in the First Place (UFP) of $x$, and thus it can be obtained directly using the $\mathtt{ilogb}(\cdot)$ function.
+For any real number $x\in\mathbb{R}$, $\lfloor\log_{2}|x|\rfloor$ represents the unit in the first place of $x$, and thus it can be obtained directly using the $\mathtt{ilogb}(\cdot)$ function.
 Following [uchino_ozaki2], the quantity $(\log_{2}(P-1)-1)/2-1$ is precomputed as $\mathcal{P}^{\prime}_{\mathit{fast}}:=\mathrm{single}(\log_{2}(P-1)/2-1.5)$.
 Furthermore, $\log_{2}\|\hat{a}_{i,:}\|_{2}$ and $\log_{2}\|\hat{b}_{:,j}\|_{2}$ can be upper-bounded by $0.51\cdot\mathtt{\_\_log2f}(\sum_{h=1}^{2k}\hat{a}_{ih}^{2})$ and $0.51\cdot\mathtt{\_\_log2f}(\sum_{h=1}^{2k}\hat{b}_{hj}^{2})$, respectively, using floating-point arithmetic in round-up mode.
-Here, $\mathtt{\_\_log2f}$ denotes the fast built-in base-2 logarithm function provided by both the CUDA Math API and the HIP math API.
-According to NVIDIA’s official CUDA documentation, the maximum absolute error of $\mathtt{\_\_log2f}(x)$ is $2^{-22}$ for $x\in[0.5,2]$, and at most 2 Units in the Last Place (ULP) otherwise [cuda].
-The definition of ULP used here follows Muller [muller:inria-00070503].
+Here, $\mathtt{\_\_log2f}$ denotes the fast built-in base-2 logarithm function provided by the CUDA Math API and the HIP math API.
+According to NVIDIA’s official CUDA documentation, the maximum absolute error of $\mathtt{\_\_log2f}(x)$ is $2^{-22}$ for $x\in[0.5,2]$, and at most 2 units in the last place (ULPs) otherwise [cuda].
+The definition of ULP used in the CUDA documentation follows Muller [muller:inria-00070503].
 In addition, the AMD HIP documentation states that the maximum absolute error of $\mathtt{\_\_log2f}(x)$ is 1 ULP for $x\in[10^{-6},10^{6}]$ [hip], although the precise definition of ULP is not specified.
-Based on these specifications, we assume the bound $\mathtt{\_\_log2f}(x)\leq(\log_{2}x)/(1-4u)$, where $u$ is the unit roundoff of FP32 arithmetic.
+Based on these specifications, we assume that the bound $\mathtt{\_\_log2f}(x)\leq(\log_{2}x)/(1-4u)$, where $u$ is the unit roundoff of FP32 arithmetic.
 If $x$ is not sufficiently close to $1$, it is first scaled to $x^{\prime}:=x/2^{\lfloor\log_{2}|x|\rfloor}$, after which the logarithm is evaluated as $\mathtt{\_\_log2f}(x^{\prime})+\lfloor\log_{2}|x|\rfloor$.
@@
 $\nu_{j}$ $:=2^{\lfloor\mathcal{P}^{\prime}_{\mathit{fast}}-\max(1,\ \delta\cdot\mathtt{\_\_log2f}(\sum_{h}\hat{b}_{hj}^{2}))\rfloor-\lfloor\log_{2}\max_{h}|\hat{b}_{hj}|\rfloor},$ (12)
-where $\mathcal{P}^{\prime}_{\mathit{fast}}-\delta\cdot x$ for $x\in\mathbb{R}$ is computed using fused multiply-add (fma) operation in round-down mode.
-In accurate mode, first, we apply diagonal scaling to $A_{R}$, $A_{I}$, $B_{R}$, and $B_{I}$ so that the integer parts of the maximum absolute values in each row of $A_{R}$ and $A_{I}$, and in each column of $B_{R}$ and $B_{I}$, fit within 6 bits.
-Specifically, we scale them as $\mathrm{diag}(\bar{\mu})A_{R}$, $\mathrm{diag}(\bar{\mu})A_{I}$, $B_{R}\mathrm{diag}(\bar{\nu})$, $B_{I}\mathrm{diag}(\bar{\nu})$, and then round them upward using the ceiling operation to obtain 7-bit or smaller integer matrices $\bar{A}_{R},\ \bar{A}_{I},\ \bar{B}_{R},\ \bar{B}_{I}$, respectively.
+where $\mathcal{P}^{\prime}_{\mathit{fast}}-\delta\cdot x$ for $x\in\mathbb{R}$ is computed using the fused multiply-add operation in round-down mode.
+In the accurate mode, we first apply diagonal scaling to $A_{R}$, $A_{I}$, $B_{R}$, and $B_{I}$ so that the integer parts of the maximum absolute values in each row of $A_{R}$ and $A_{I}$ and in each column of $B_{R}$ and $B_{I}$ fit within 6 bits.
+Specifically, we scale them as $\mathrm{diag}(\bar{\mu})A_{R}$, $\mathrm{diag}(\bar{\mu})A_{I}$, $B_{R}\mathrm{diag}(\bar{\nu})$, and $B_{I}\mathrm{diag}(\bar{\nu})$, and then round them up using the ceiling operation to obtain 7-bit or smaller integer matrices $\bar{A}_{R}$, $\bar{A}_{I}$, $\bar{B}_{R}$, and $\bar{B}_{I}$, respectively.
 These matrices satisfy
@@
 where no computing error occurs in $\bar{A}_{R}-\bar{A}_{I}$ and $\bar{B}_{R}-\bar{B}_{I}$ in INT8.
-Then, the elements of $\mu$ and $\nu$ can be determined in the same manner as in the real-valued Ozaki-II scheme [uchino_ozaki2], i.e.,
+Then, the elements of $\mu$ and $\nu$ can be determined in the same manner as in the real-valued Ozaki-II scheme [uchino_ozaki2]; that is,
 $\mu_{i}$ $:=\bar{\mu}_{i}\cdot 2^{\lfloor\mathcal{P}^{\prime}_{\mathit{accu}}-\delta\cdot\log_{2}(\max_{h}\max((\bar{C}_{R})_{ih},(\bar{C}_{I})_{ih}))\rfloor},$ (13)
 $\nu_{j}$ $:=\bar{\nu}_{j}\cdot 2^{\lfloor\mathcal{P}^{\prime}_{\mathit{accu}}-\delta\cdot\log_{2}(\max_{h}\max((\bar{C}_{R})_{hj},(\bar{C}_{I})_{hj}))\rfloor},$ (14)
-where $\mathcal{P}^{\prime}_{\mathit{accu}}:=\mathrm{single}(\log_{2}(P-1)/2-0.5)$, and $\mathcal{P}^{\prime}_{\mathit{accu}}-\delta\cdot x$ for $x\in\mathbb{R}$ is computed using fma in round-down mode.
+where $\mathcal{P}^{\prime}_{\mathit{accu}}:=\mathrm{single}(\log_{2}(P-1)/2-0.5)$ and $\mathcal{P}^{\prime}_{\mathit{accu}}-\delta\cdot x$ for $x\in\mathbb{R}$ is computed using the fused multiply-add operation in round-down mode.
 III-C Performance Analysis
-To analyze the performance characteristics of the proposed method, we employ a simple yet expressive performance model that captures both computation and memory-transfer costs.
-Let $b$ denote the peak sustained memory bandwidth (in Byte/s) and $p$ denote the peak throughput of the INT8 matrix-multiplication engine (in OPS).
-For a given operation, the memory-bound execution time is estimated as the required data movement divided by $b$, while the compute-bound time is obtained as the arithmetic workload divided by $p$.
+To analyze the performance characteristics of the proposed methods, we employ a simple yet expressive performance model that captures both computation and memory transfer costs.
+Let $b$ denote the peak sustained memory bandwidth (in B/s) and $p$ denote the peak throughput of the INT8 matrix-multiplication engine (in operations per second, OPS).
+For a given operation, the memory-bound execution time is estimated as the required data movement divided by $b$ and the compute-bound time is obtained as the arithmetic workload divided by $p$.
 Because even memory-bound kernels incur a non-negligible amount of arithmetic overhead, we include a correction term $c$.
 III-C1 Fast Mode
-Let the input matrices $A\in\mathbb{C}^{m\times k}$ and $B\in\mathbb{C}^{k\times n}$ be double precision.
-In step 1, both $A$ and $B$ are loaded twice: once for computing the rowwise maximum absolute values of $\hat{A}$ and the columnwise maximum absolute values of $\hat{B}$, and once for computing the residues of $\hat{A}$ and $\hat{B}$.
+Let the input matrices $A\in\mathbb{C}^{m\times k}$ and $B\in\mathbb{C}^{k\times n}$ be double-precision.
+In step 1, both $A$ and $B$ are loaded twice: once for computing the row-wise maximum absolute values of $\hat{A}$ and the column-wise maximum absolute values of $\hat{B}$, and once for computing the residues of $\hat{A}$ and $\hat{B}$.
 Afterward, the scaling vectors and the INT8 matrices are stored.
 Each scaling factor is a power of two, so only its exponent needs to be stored as an INT16.
-Including the correction term $c$, which accounts for arithmetic overhead even in memory-bound operations, the total execution time of step 1 can be estimated as $((32+3N+c)k+2)(m+n)/b$.
+Including the correction term $c$, which accounts for arithmetic overhead even in memory-bound operations, the total execution time for step 1 can be estimated as $((32+3N+c)k+2)(m+n)/b$.
 In step 2, the INT8 GEMM kernels are executed $3N$ times.
 The resulting $3N$ INT32 matrices are then loaded and converted into $N$ INT8 complex matrices.
-Accounting for computational overhead $c$, the execution time of step 2 is estimated as $6Nmnk/p+(14N+c)mn/b$.
-In step 3, the $N$ INT8 complex matrices are loaded, and the CRT accumulation and final reduction are performed.
+Accounting for computational overhead $c$, the execution time for step 2 is estimated as $6Nmnk/p+(14N+c)mn/b$.
+In step 3, the $N$ INT8 complex matrices are loaded and the CRT accumulation and final reduction are performed.
 The scaling vectors are then loaded, the inverse scaling is applied, and finally the resulting double-precision complex matrix is stored.
-With the correction term $c$, the execution time of step 3 is estimated as $(2(m+n)+(2N+16+c)mn)/b$.
-Combining the costs derived for steps 1, 2, and 3, the total execution time of the proposed method can be estimated as the sum of its memory-bound and compute-bound components.
+With the correction term $c$, the execution time for step 3 is estimated as $(2(m+n)+(2N+16+c)mn)/b$.
+Combining the costs derived for steps 1, 2, and 3, the total execution time for the proposed methods can be estimated as the sum of its memory-bound and compute-bound components.
 Aggregating all terms yields the following expression for the overall execution time:
@@
 $\begin{multlined}\frac{((35+3N+c)k+8)(m+n)+(16N+40+2c)mn}{b}\\ +\frac{6(N+1)mnk}{p},\end{multlined}\frac${((35+3N+c)k+8)(m+n)+(16N+40+2c)mn}{b}\\ $\frac${6(N+1)mnk}{p},
-and for the single-precision case,
+and for the single-precision case as
 $\begin{multlined}\frac{((19+3N+c)k+8)(m+n)+(16N+32+2c)mn}{b}\\ +\frac{6(N+1)mnk}{p}.\end{multlined}\frac${((19+3N+c)k+8)(m+n)+(16N+32+2c)mn}{b}\\ $\frac${6(N+1)mnk}{p}.
-Figs. 2 and 3 show performance-model heatmaps for the fast and accurate modes in single and double precision for $6$ and $13$ moduli, respectively. The TFLOPS was computed as
+Figs. 2 and 3 show performance model heatmaps for the fast and accurate modes in single and double precision for $6$ and $13$ moduli, respectively. The TFLOPS was computed as
 $\frac{8mnk}{t}\cdot 10^{-12},$
 where $t$ is the execution time in seconds.
-For example, assuming an effective memory bandwidth of 2–4 TByte/s and an observed INT8 GEMM kernel throughput of 1500 TFLOPS on the GH200, the model predicts that ZGEMM emulation in the accurate mode would achieve approximately 120 TFLOPS.
-Figure 2: Performance-model heatmaps for the single-precision complex matrix multiplication emulation.
-The left panel corresponds to the fast mode and the right panel to the accurate mode.
-The horizontal axis denotes the attainable memory bandwidth, and the vertical axis denotes the achievable INT8 GEMM throughput.
-The problem size is fixed to $m=n=k=16384$, and the correction term is set to $c=6$, equal to the number of moduli used in the emulation.
-The color scale indicates the predicted throughput in TFLOPS of the proposed emulation.
-Figure 3: Performance-model heatmaps for the double-precision complex matrix multiplication emulation.
-The left panel corresponds to the fast mode and the right panel to the accurate mode.
-The horizontal axis denotes the attainable memory bandwidth, and the vertical axis denotes the achievable INT8 GEMM throughput.
-The problem size is fixed to $m=n=k=16384$, and the correction term is set to $c=13$, equal to the number of moduli used in the emulation.
-The color scale indicates the predicted throughput in TFLOPS of the proposed emulation.
+For example, assuming an effective memory bandwidth of 2–4 TB/s and an observed INT8 GEMM kernel throughput of 1500 TFLOPS on the GH200, the model predicts that ZGEMM emulation in the accurate mode would achieve approximately 120 TFLOPS.
+Figure 2: Performance model heatmaps for single-precision complex matrix multiplication emulation.
+The left panel corresponds to the fast mode and the right panel corresponds to the accurate mode.
+The horizontal axis denotes the achievable memory bandwidth and the vertical axis denotes the achievable INT8 GEMM throughput.
+The problem size is fixed at $m=n=k=16384$ and the correction term is set to $c=6$, equal to the number of moduli used in the emulation.
+The color scale indicates the predicted throughput (in TFLOPS) of the proposed emulation.
+Figure 3: Performance model heatmaps for double-precision complex matrix multiplication emulation.
+The left panel corresponds to the fast mode and the right panel corresponds to the accurate mode.
+The horizontal axis denotes the achievable memory bandwidth and the vertical axis denotes the achievable INT8 GEMM throughput.
+The problem size is fixed at $m=n=k=16384$ and the correction term is set to $c=13$, equal to the number of moduli used in the emulation.
+The color scale indicates the predicted throughput (in TFLOPS) of the proposed emulation.
 III-D GEMM Emulation Library
@@
 Details of the library and usage instructions are provided in [GEMMul8].
-The library supports both CUDA and HIP backends, and the CUDA-based source code is automatically translated into HIP C++ when compiled under the HIP backend.
-The implementation provides emulation routines for SGEMM and DGEMM as proposed in previous work, as well as the CGEMM and ZGEMM emulation kernels developed in the present study.
-The supported data types are float, double, cuComplex, cuDoubleComplex, hipFloatComplex, hipDoubleComplex.
-To enable transparent integration with existing HPC applications, the library can be activated by adding the path of the shared library to the environment variable LD_PRELOAD.
+The library supports both CUDA and HIP backends.
+The CUDA-based source code is automatically translated into HIP C++ when compiled under the HIP backend.
+The implementation provides emulation routines for single- and double-precision GEMM, proposed in previous work [uchino_ozaki2], as well as the CGEMM and ZGEMM emulation kernels developed in the present study.
+The supported data types are float, double, cuComplex, cuDoubleComplex, hipFloatComplex, and hipDoubleComplex.
+To enable transparent integration with existing high-performance computing applications, the library can be activated by adding the path of the shared library to the environment variable LD_PRELOAD.
 Once activated, the following GEMM calls to CUDA or HIP backends are dynamically redirected to the emulation routines, allowing users to evaluate GEMM emulation without modifying their applications:
@@
   • hipblasGemmEx, hipblasGemmEx_v2
-The implementation consists of INT8 matrix multiplications, fixed-order basic arithmetic operations, and the evaluation of the $\log_{2}(\cdot)$ function appearing in (11), (12), (13), and (14).
+The implementation consists of INT8 matrix multiplications, fixed-order basic arithmetic operations, and the evaluation of the $\log_{2}(\cdot)$ function in (11), (12), (13), and (14).
 Rounding errors arise only in the computation of the scaling vectors defined by (11), (12), (13), and (14), and in the accumulation and final reduction stages of the CRT.
@@
 IV Numerical Results
-Here, we compare the following methods with each other:
+Here, we compare the following methods:
 Cgemm:
@@
 Ozaki-I from cuBLAS in Fixed Mantissa Control mode with $S$ slices
-The hipBLAS does not provide hipblasCgemm3m or hipblasZgemm3m and the same limitation applies to the rocBLAS, though it is not used in this study.
-We have empirically confirmed that both the performance and numerical accuracy of cublasCgemm3m are essentially identical to those of cublasCgemm; thus, we omit its results from the following comparisons.
+hipBLAS does not provide hipblasCgemm3m or hipblasZgemm3m; the same limitation applies to rocBLAS, though it is not used in this study.
+We empirically confirmed that both the performance and numerical accuracy of cublasCgemm3m are essentially identical to those of cublasCgemm; thus, we omit the former’s results from the following comparisons.
 The cuBLAS implementation of Ozaki-I-based emulation provides two execution modes: Dynamic Mantissa Control and Fixed Mantissa Control.
-In the Dynamic Mantissa Control mode, the routine determines the number of mantissa bits, equivalently, the number of slices, required for fixed-point emulation such that the numerical accuracy matches or exceeds that of FP64.
+In the Dynamic Mantissa Control mode, the routine determines the number of mantissa bits, or equivalently the number of slices, required for fixed-point emulation such that the numerical accuracy matches or exceeds that of FP64.
 If the required mantissa length exceeds either an internal limit or a user-specified maximum, the computation is automatically executed in native FP64.
-In the Fixed Mantissa Control mode, the user specifies the mantissa-bit length used for the fixed-point representation.
+In the Fixed Mantissa Control mode, the user specifies the mantissa bit length used for the fixed-point representation.
 This mode reduces runtime overhead by eliminating dynamic precision selection, but it does not ensure FP64-level accuracy.
@@
 where $\mathrm{rand}\in(0,1]$ denotes a uniform random number and $\mathrm{randn}$ denotes a standard normal random number.
-Both types of random numbers were generated using either the cuRAND API or the hipRAND library.
-Following the methodology adopted in prior studies about the Ozaki schemes [ozaki2012error, ozaki2013generalization, mukunoki2020, ootomo2024dgemm, uchino2025Performance, ozaki-scheme2, uchino_ozaki2], we introduce the parameter $\phi$ to control the dynamic range of the input matrices.
+Both types of random number were generated using either the cuRAND API or the hipRAND library.
+Following the methodology adopted in prior studies on Ozaki schemes [ozaki2012error, ozaki2013generalization, mukunoki2020, ootomo2024dgemm, uchino2025Performance, ozaki-scheme2, uchino_ozaki2], we introduce the parameter $\phi$ to control the dynamic range of the input matrices.
 In accordance with these studies, we specify $\phi\in\{0,0.5,1,1.5\}$ for single-precision tests and $\phi\in\{0.5,1,2,4\}$ for double-precision tests.
@@
 This confirms that $\phi$ effectively widens the dynamic range without substantially affecting the central tendency of the distribution.
-Such controlled variation is essential for evaluating the performance and numerical stability of the Ozaki schemes.
-Figs. 4 and 5 show the accuracy of CGEMM and ZGEMM, respectively, observed on the GH200. Similar results were obtained on other GPUs. The maximum relative error in the figures were calculated as
+Such controlled variation is essential for evaluating the performance and numerical stability of Ozaki schemes.
+Figs. 4 and 5 show the accuracy of CGEMM and ZGEMM, respectively, observed on the GH200. Similar results were obtained on other GPUs. The maximum relative error in the figures was calculated as
 $\max_{i,j}\left(\max\left(\frac{|\mathrm{Re}(\tilde{c}_{ij})-\mathrm{Re}(c_{ij})|}{|\mathrm{Re}(c_{ij})|},\frac{|\mathrm{Im}(\tilde{c}_{ij})-\mathrm{Im}(c_{ij})|}{|\mathrm{Im}(c_{ij})|}\right)\right),$
 where $\tilde{C}\in\mathbb{C}^{m\times n}$ is the approximation of $AB$ obtained using the emulation methods and the standard complex GEMM routines, and $C\in\mathbb{C}^{m\times n}$ is the high-precision result of $AB$ computed using double-double arithmetic.
-In CGEMM emulation, fast-$N$ with $6\leq N\leq 9$ and accu-$N$ with $6\leq N\leq 8$ provided the accuracy comparable with that of CGEMM.
-In ZGEMM emulation, fast-$N$ with $13\leq N\leq 18$ and accu-$N$ with $13\leq N\leq 17$ achieved the accuracy comparable with that of ZGEMM.
-As $\phi$ increases, the range of exponents for $A$ and $B$ expands, making it easier to overestimate the upper bound $2\sum_{h=1}^{k}|a^{\prime}_{ih}||b^{\prime}_{hj}|$ in the condition (4).
-DGEMM emulation in [uchino_ozaki2] required at least 14 moduli to achieve DGEMM-level accuracy, whereas the proposed method attains ZGEMM-level accuracy with only 13 moduli.
+In CGEMM emulation, fast-$N$ with $6\leq N\leq 9$ and accu-$N$ with $6\leq N\leq 8$ provided accuracy comparable to that of the native CGEMM.
+In ZGEMM emulation, fast-$N$ with $13\leq N\leq 18$ and accu-$N$ with $13\leq N\leq 17$ achieved accuracy comparable to that of the native ZGEMM.
+As $\phi$ increases, the range of exponents for $A$ and $B$ expands, making it easier to overestimate the upper bound $2\sum_{h=1}^{k}|a^{\prime}_{ih}||b^{\prime}_{hj}|$ in condition (4).
+DGEMM emulation in [uchino_ozaki2] required at least 14 moduli to achieve native DGEMM-level accuracy, whereas the proposed method attains native ZGEMM-level accuracy with only 13 moduli.
 This reduction is attributed to the use of the Karatsuba algorithm in the INT8 complex matrix multiplication, which suppresses precision loss more effectively than floating-point arithmetic.
@@
 Enhancing the limiting accuracy therefore necessitates performing these operations with even higher precision.
-Figure 4: Maximum relative error of single-precision complex matrix multiplication on NVIDIA GH200 Grace Hopper Superchip, CUDA Toolkit 13.0.88, and gcc 11.5.0, for $m=n=1024$ and $k=16384$.
-Figure 5: Maximum relative error of double-precision complex matrix multiplication on NVIDIA GH200 Grace Hopper Superchip, CUDA Toolkit 13.0.88, and gcc 11.5.0, for $m=n=1024$ and $k=16384$.
+Figure 4: Maximum relative error of single-precision complex matrix multiplication on NVIDIA GH200 Grace Hopper Superchip (CUDA Toolkit 13.0.88, gcc 11.5.0) for $m=n=1024$ and $k=16384$.
+Figure 5: Maximum relative error of double-precision complex matrix multiplication on NVIDIA GH200 Grace Hopper Superchip (CUDA Toolkit 13.0.88, gcc 11.5.0) for $m=n=1024$ and $k=16384$.
 IV-B Performance
 […]
-Figure 6: Throughput performance in TFLOPS of single-precision complex matrix multiplication on NVIDIA GH200 Grace Hopper Superchip, CUDA Toolkit 13.0.88, and gcc 11.5.0, for $m=k=n$.
-Methods shown in red italic achieve CGEMM-level accuracy.
-Figure 7: Throughput performance in TFLOPS of single-precision complex matrix multiplication on NVIDIA B200 Blackwell GPU with Intel Xeon Platinum 8570 CPU, CUDA Toolkit 12.8.93, and g++ 13.3.0, for $m=k=n$.
-Methods shown in red italic achieve CGEMM-level accuracy.
-Figure 8: Throughput performance in TFLOPS of single-precision complex matrix multiplication on NVIDIA GeForce RTX 5080 GPU with AMD Ryzen 7 3700X CPU, CUDA Toolkit 13.0.88, and gcc 11.5.0, for $m=k=n$.
-Methods shown in red italic achieve CGEMM-level accuracy.
-Figure 9: Throughput performance in TFLOPS of single-precision complex matrix multiplication on AMD Instinct MI300X GPU with AMD EPYC 9534 CPU, ROCm 6.4.2, and clang 19.0.0git, for $m=k=n$.
-Methods shown in red italic achieve CGEMM-level accuracy.
-Figure 10: Throughput performance in TFLOPS of double-precision complex matrix multiplication on NVIDIA GH200 Grace Hopper Superchip, CUDA Toolkit 13.0.88, and gcc 11.5.0, for $m=k=n$.
-Methods shown in red italic achieve ZGEMM-level accuracy.
-Figure 11: Throughput performance in TFLOPS of double-precision complex matrix multiplication on NVIDIA B200 Blackwell GPU with Intel Xeon Platinum 8570 CPU, CUDA Toolkit 12.8.93, and g++ 13.3.0, for $m=k=n$.
-Methods shown in red italic achieve ZGEMM-level accuracy.
-Figure 12: Throughput performance in TFLOPS of double-precision complex matrix multiplication on NVIDIA GeForce RTX 5080 GPU with AMD Ryzen 7 3700X CPU, CUDA Toolkit 13.0.88, and gcc 11.5.0, for $m=k=n$.
-Methods shown in red italic achieve ZGEMM-level accuracy.
-Figure 13: Throughput performance in TFLOPS of double-precision complex matrix multiplication on AMD Instinct MI300X GPU with AMD EPYC 9534 CPU, ROCm 6.4.2, and clang 19.0.0git, for $m=k=n$.
-Methods shown in red italic achieve ZGEMM-level accuracy.
+Figure 6: Throughput performance (in TFLOPS) of single-precision complex matrix multiplication on NVIDIA GH200 Grace Hopper Superchip (CUDA Toolkit 13.0.88, gcc 11.5.0) for $m=k=n$.
+Methods shown in red and italics achieve CGEMM-level accuracy.
+Figure 7: Throughput performance (in TFLOPS) of single-precision complex matrix multiplication on NVIDIA B200 Blackwell GPU with Intel Xeon Platinum 8570 CPU (CUDA Toolkit 12.8.93, g++ 13.3.0) for $m=k=n$.
+Methods shown in red and italics achieve CGEMM-level accuracy.
+Figure 8: Throughput performance (in TFLOPS) of single-precision complex matrix multiplication on NVIDIA GeForce RTX 5080 GPU with AMD Ryzen 7 3700X CPU (CUDA Toolkit 13.0.88, gcc 11.5.0) for $m=k=n$.
+Methods shown in red and italics achieve CGEMM-level accuracy.
+Figure 9: Throughput performance (in TFLOPS) of single-precision complex matrix multiplication on AMD Instinct MI300X GPU with AMD EPYC 9534 CPU (ROCm 6.4.2, clang 19.0.0git) for $m=k=n$.
+Methods shown in red and italics achieve CGEMM-level accuracy.
+Figure 10: Throughput performance (in TFLOPS) of double-precision complex matrix multiplication on NVIDIA GH200 Grace Hopper Superchip (CUDA Toolkit 13.0.88, gcc 11.5.0) for $m=k=n$.
+Methods shown in red and italics achieve ZGEMM-level accuracy.
+Figure 11: Throughput performance (in TFLOPS) of double-precision complex matrix multiplication on NVIDIA B200 Blackwell GPU with Intel Xeon Platinum 8570 CPU (CUDA Toolkit 12.8.93, g++ 13.3.0) for $m=k=n$.
+Methods shown in red and italics achieve ZGEMM-level accuracy.
+Figure 12: Throughput performance (in TFLOPS) of double-precision complex matrix multiplication on NVIDIA GeForce RTX 5080 GPU with AMD Ryzen 7 3700X CPU (CUDA Toolkit 13.0.88, gcc 11.5.0) for $m=k=n$.
+Methods shown in red and italics achieve ZGEMM-level accuracy.
+Figure 13: Throughput performance (in TFLOPS) of double-precision complex matrix multiplication on AMD Instinct MI300X GPU with AMD EPYC 9534 CPU (ROCm 6.4.2, clang 19.0.0git) for $m=k=n$.
+Methods shown in red and italics achieve ZGEMM-level accuracy.
+Figure 14: Speedups of single-precision complex matrix multiplication over native CGEMM for $m=k=n=16384$
+Figure 15: Speedups of double-precision complex matrix multiplication over native ZGEMM for $m=k=n=16384$
 These performance characteristics correlate strongly with the architectural balance of arithmetic throughput on each GPU.
-The GH200 provides a ratio of $\text{FP64}:\text{FP32}:\text{INT8}=1:1:29.5$, indicating that INT8 operations are moderately more abundant than FP64 and FP32 units.
+The GH200 provides a ratio of $\text{FP64}:\text{FP32}:\text{INT8}=1:1:29.5$, indicating that INT8 units are more abundant than FP64 and FP32 units.
 The B200 exhibits an even more pronounced imbalance, $1:2:121.6$, where INT8 throughput overwhelmingly dominates the floating-point pipelines.
 This architectural skew directly amplifies the benefit of the proposed kernels, which replace large portions of floating-point arithmetic with INT8-intensive computation.
-In contrast, the MI300X offers a more subdued ratio of $1:1:16$, providing substantially less INT8 advantage relative to its floating-point throughput.
+In contrast, the MI300X offers a more moderate ratio of $1:1:16$, with INT8 throughput being moderately higher than floating-point throughput.
 As a result, the computational substitution enabled by our method yields only marginal acceleration on the MI300X, consistent with the limited speedups observed in the benchmarks.
-For the RTX 5080, the observed performance gains, especially the very large speedups in ZGEMM, suggest a much stronger favoring of INT8 throughput relative to floating-point units.
-The architectural balance on RTX 5080 corresponds to $\text{FP64}:\text{FP32}:\text{INT8}=1:64:512$, which explains the pronounced acceleration when using INT8-based emulation.
-Compared with the Ozaki-I emulation implemented in cuBLAS, the proposed method delivered substantially higher throughput on both GH200 and RTX 5080 GPUs.
-On the GH200, Ozaki-I with 7–9 slices achieved 25–39 TFLOPS, whereas the proposed method attained 99–136 TFLOPS.
-On the RTX 5080, Ozaki-I achieved 6.4–10 TFLOPS, while the proposed method reached 19–26 TFLOPS.
-These disparities arise from the fundamental algorithmic differences: Ozaki-I with $S$ slices requires $S(S+1)/2$ INT8 complex matrix multiplications, whereas the proposed technique with $N$ moduli requires only $N$ INT8 complex matrix multiplications and and further reduces their cost by a factor of $3/4$ using the Karatsuba method.
-This reduction in both the number and cost of INT8 matrix multiplication kernel invocations directly explains the substantially higher performance of the proposed method.
-IV-C Supplemental Comparison for Real-Valued Emulation
-Here, we focus on double-precision real matrix multiplication.
+For the RTX 5080, the observed performance gains, especially the very large speedups in ZGEMM, suggest a much higher INT8 throughput relative to that of floating-point units.
+The architectural balance on the RTX 5080 corresponds to $\text{FP64}:\text{FP32}:\text{INT8}=1:64:512$, which explains the pronounced acceleration when using INT8-based emulation.
+Compared to the Ozaki-I emulation implemented in cuBLAS, the proposed methods delivered substantially higher throughput on both the GH200 and RTX 5080.
+On the GH200, Ozaki-I with 7–9 slices achieved 25–39 TFLOPS, whereas the proposed methods achieved 99–136 TFLOPS.
+On the RTX 5080, Ozaki-I achieved 6.4–10 TFLOPS, whereas the proposed methods achieved 19–26 TFLOPS.
+These disparities arise from fundamental algorithmic differences: Ozaki-I with $S$ slices requires $S(S+1)/2$ INT8 complex matrix multiplications, whereas the proposed technique with $N$ moduli requires only $N$ INT8 complex matrix multiplications, whose cost is further reduced by a factor of $3/4$ using the Karatsuba method.
+This reduction in both the number and cost of INT8 matrix multiplication kernel invocations explains the substantially higher performance of the proposed methods.
+IV-C Supplemental Comparison of Real-Valued Emulation
 In this section, we focus on double-precision real matrix multiplication.
@@
 According to [uchino_ozaki2], the fast mode and the accurate mode require 14–18 and 14–17 moduli, respectively, to achieve DGEMM-level accuracy.
-On the GH200 with matrix size $m=n=k=16384$, the Ozaki-II fast-$N$ configuration with $14\leq N\leq 18$ achieved 63–82 TFLOPS, while the accu-$N$ configuration with $14\leq N\leq 17$ reached 63–75 TFLOPS.
-Incorporating the blocking technique improved performance across all settings: the blocked fast-$N$ version attained 72–93 TFLOPS, and the blocked accu-$N$ version delivered 71–85 TFLOPS.
-cuBLAS’s Ozaki-I implementation with 7–10 slices achieved only 20–39 TFLOPS, and the standard DGEMM kernel sustained 61 TFLOPS.
-On the RTX 5080 with $m=n=k=8192$, the fast-$N$ and accu-$N$ configurations of the proposed method reached 12–16 TFLOPS and 12–15 TFLOPS, respectively, while cuBLAS’s Ozaki-I achieved 5.4–10 TFLOPS.
-In comparison, the native DGEMM kernel delivered 0.84 TFLOPS, highlighting the substantial advantage of INT8-intensive Ozaki-type emulation on this consumer architecture.
-These results demonstrate that blocking substantially enhances the performance of Ozaki-II for large-scale problems, particularly on bandwidth-rich GPUs such as the H200, and Ozaki-II, whether blocked or unblocked, consistently outperforms cuBLAS’s Ozaki-I due to its reduced number of INT8 matrix multiplication invocations.
+On the GH200 with matrix size $m=n=k=16384$, the Ozaki-II fast-$N$ configuration with $14\leq N\leq 18$ achieved 63–82 TFLOPS and the accu-$N$ configuration with $14\leq N\leq 17$ achieved 63–75 TFLOPS.
+Incorporating the blocking technique improved performance across all settings: the blocked fast-$N$ version achieved 72–93 TFLOPS and the blocked accu-$N$ version achieved 71–85 TFLOPS.
+cuBLAS’s Ozaki-I implementation with 7–10 slices achieved only 20–39 TFLOPS and the standard DGEMM kernel achieved 61 TFLOPS.
+On the RTX 5080 with $m=n=k=8192$, the fast-$N$ and accu-$N$ configurations of the proposed methods achieved 12–16 TFLOPS and 12–15 TFLOPS, respectively, and cuBLAS’s Ozaki-I achieved 5.4–10 TFLOPS.
+In comparison, the native DGEMM kernel achieved 0.84 TFLOPS, highlighting the substantial advantage of INT8-intensive Ozaki-type emulation on this consumer architecture.
+These results demonstrate that blocking substantially enhances the performance of Ozaki-II for large-scale problems, particularly on bandwidth-rich GPUs such as the B200, and that Ozaki-II, whether blocked or unblocked, consistently outperforms cuBLAS’s Ozaki-I due to its reduced number of INT8 matrix multiplication invocations.
 V Conclusion
-This study proposed emulation methods of complex matrix multiplication based on the Ozaki-II scheme and developed a portable library that supports execution on both AMD and NVIDIA GPUs.
+This study proposed emulation methods for complex matrix multiplication based on the Ozaki-II scheme and developed a portable library that supports execution on both AMD and NVIDIA GPUs.
 The numerical results demonstrated that the proposed methods outperform the standard GEMM routines in cuBLAS and hipBLAS, which are implemented using native FP32 and FP64 arithmetic.
-In addition, the proposed methods outperform the emulation using the Ozaki-I scheme implemented in cuBLAS.
-Future work includes comparisons with the BF16x9 CGEMM emulation algorithm provided for NVIDIA B200 and B300 GPUs in cuBLAS 12.9 and later.
-A major limitation of the proposed method, as well as emulation-based approaches in general, is the substantial working memory required.
+In addition, the proposed methods outperform emulation using the Ozaki-I scheme implemented in cuBLAS.
+Future work will include comparisons with the BF16x9 CGEMM emulation algorithm provided for NVIDIA B200 and B300 GPUs introduced in cuBLAS 12.9.
+A major limitation of the proposed methods, as well as emulation-based approaches in general, is the substantial working memory required.
 This overhead is currently unavoidable when aiming to achieve high performance without relying on FP32 or FP64 matrix multiplication.
-Addressing this issue remains an open challenge not only for emulation techniques but also for HPC applications that employ them.
-The Ozaki-I and Ozaki-II schemes utilize INT8 matrix engines, benefiting from the fact that the error-free low-precision matrix multiplications appearing in these schemes do not require exponent handling, enabling efficient synergy with INT8 operations with INT32 accumulation.
-When using FP16 or BF16 matrix engines with FP32 accumulation, the effective representable integer range is limited to 24 bits due to the FP32 significand, leading to significantly reduced computational efficiency compared with INT32 accumulation.
+Addressing this issue remains an open challenge not only for emulation techniques but also for high-performance computing applications that employ them.
+The Ozaki-I and Ozaki-II schemes utilize INT8 matrix engines, benefiting from the fact that the error-free low-precision matrix multiplications in these schemes do not require exponent handling, enabling efficient synergy with INT8 operations with INT32 accumulation.
+When using FP16 or BF16 matrix engines with FP32 accumulation, the effective representable integer range is limited to 24 bits due to the FP32 significand, leading to significantly reduced computational efficiency compared to that of INT32 accumulation.
 However, some recent GPUs, such as the NVIDIA B300, provide reduced INT8 capabilities.
-Supporting such hardware architectures represents an additional direction for future work.
+Support for such hardware architectures will be considered in future work.
 Acknowledgment